### v2

# Is it Really Negative? Evaluating Natural Language Video Localization Performance on Multiple Reliable Videos Pool

Abstract
With the explosion of multimedia content in recent years, Video Corpus Moment Retrieval (VCMR), which aims to detect a video moment that matches a given natural language query from multiple videos, has become a critical problem. However, existing VCMR studies have a significant limitation since they have regarded all videos not paired with a specific query as negative, neglecting the possibility of including false negatives when constructing the negative video set. In this paper, we propose an MVMR (Massive Videos Moment Retrieval) task that aims to localize video frames within a massive video set, mitigating the possibility of falsely distinguishing positive and negative videos. For this task, we suggest an automatic dataset construction framework by employing textual and visual semantic matching evaluation methods on the existing video moment search datasets and introduce three MVMR datasets. To solve MVMR task, we further propose a strong method, CroCs, which employs cross-directional contrastive learning that selectively identifies the reliable and informative negatives, enhancing the robustness of a model on MVMR task. Experimental results on the introduced datasets reveal that existing video moment search models are easily distracted by negative video frames, whereas our model shows significant performance. [^1] We have opened an anonymous GitHub page to make our code and dataset publicly available: link
1 Introduction
Enabled by the increased accessibility of video-sharing platforms along with the advancement of networking and storage technology, a vast number of new content is available on the web daily. With this huge number of contents, Natural Language Video Localization (NLVL) task, searching for a video moment suitable for a given query, has emerged as an essential problem (Anne Hendricks et al., 2017; Gao et al., 2017). However, this line of research narrows down the task to an impractical setting by targeting the search over a single video; thus, it bears limitations in retrieving relevant information over whole media pools.
Figure 1: MVMR vs. VCMR. In existing VCMR studies, only a single golden positive video moment is identified as positive, while entire videos not paired with a specific query are regarded as negative. Our proposed MVMR filters positive video moments for a query from the whole video set; thus, it covers more practical and generalized use cases with reliable settings. $v^{+}_{i}$ and $v^{-}_{j}$ mean a positive and a negative video, respectively.
Therefore, Video Corpus Moment Retrieval (VCMR) task that retrieves a correct positive moment among numerous negative videos has emerged (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022). Nonetheless, VCMR also has significant limitations as it merely extends NLVL datasets by categorizing all unpaired moments for a specific text query as negative. This assumption diverges from the practical setting of video moment search where multiple positive videos exist in whole media collections. Furthermore, existing VCMR studies have overlooked the potential presence of false-negative videos within the unpaired video set when constructing video retrieval pools for a specific query. As a result, evaluating the performance of NLVL models using an existing VCMR framework may lead to unreliable results. For example, in TACoS test dataset, it is noteworthy that the number of videos with queries exactly matching the text "The person gets out a knife." amounts to 60% of all the videos. It is a significant problem since more videos will likely be incorrectly classified as negatives when expanding the matching criteria to include broader semantic similarities.
To bridge these gaps, we propose an MVMR (Massive Videos Moment Retrieval) task, which expands the search coverage to a massive video set that can include any number of positive videos as shown in Figure 1. Specifically, given a query $q$ with a positive videos $v^{+}_{1},...,v^{+}_{k}$ and a negative videos $v^{-}_{1},...,v^{-}_{n-k}$, the MVMR task aims to detect temporal moments in the positive videos that matches the query from a massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$. Furthermore, we propose simple and effective methods to construct a practical MVMR dataset by leveraging publicly available NLVL datasets. A critical challenge in converting an NLVL dataset into an MVMR dataset is accurately defining positive and negative videos for each query. However, manually performing this task by humans is exceedingly labor-intensive due to the massive volume of videos that need to be evaluated for each query. To address this issue, we propose an automated framework for MVMR dataset construction. This framework employs textual and visual semantic matching evaluation methods to define MVMR positive and negative videos while mitigating the danger of false categorization. Specifically, we examine the similarity between a target query and all videos in an NLVL dataset using SimCSE Gao et al. (2021) and EMScore Shi et al. (2022) models to define a massive video set. To this end, we construct three MVMR datasets, and we confirm that three samples of the introduced datasets contain only 1.5%, 0.2%, and 3.5% falsely defined videos, showing that our approach successfully creates a practical benchmark dataset from human evaluation.
To tackle MVMR task, we also introduce a novel negative sample training method, Cross-directional Hard and Reliable Negative Contrastive Learning (CroCs). Specifically, CroCs adopts two training mechanisms: (1) weakly-supervised potential negative learning by assessing the similarity between queries, and (2) hard-negative learning by utilizing the relevance score between queries and moments predicted by a model, which helps find challenging negative samples. We demonstrate the performance of strong baseline models, including a state-of-the-art model, to compare with our model in the new scenario (MVMR). The experimental results show that the performance of baseline models significantly degrades on the new task. For example, the state-of-the-art model, MMN, shows average 29.7% degradation on R@1(IOU0.5) for all MVMR datasets, revealing the difficulty of the MVMR task. In contrast, our CroCs outperforms baseline models significantly in the MVMR setting, and this result proves that informative negative samples enhance the performance of the MVMR task. Moreover, we solve the MVMR task in a more practical scenario by adopting a pipeline with a video retrieval model, revealing that CroCs also outperforms the strong baseline.
2 Backgrounds
2.1 Natural Language Video Localization
Natural language video localization (NLVL), which aims to detect a specific moment in a video that matches a given text query, is one of the most challenging problems in the video-language multi-modal domain (Zhang et al., 2020b, a; Nan et al., 2021; Gao and Xu, 2021; Liu et al., 2021; Wang et al., 2022). Therefore, the NLVL task has been explored extensively, with numerous effective methods emerging. Some novel approaches have regarded the problem as question-answering (QA) tasks (Seo et al., 2018; Joshi et al., 2020), by leveraging a QA model to encode a multi-modal representation and then predicting the frames that correspond to the start and end of the moment that match a query (Ghosh et al., 2019; Zhang et al., 2020a). However, NLVL studies operate within an impractical framework, focusing on searching within a single video. Consequently, they have limitations in retrieving relevant information across the whole media repositories.
     Dataset   # Queries / # Videos       Avg Len (sec)       Avg # Moments       Avg      # Max Positive /
Moment / Video per Query Query Len # Retrieval Pool
Charades-STA 3720 / 1334 7.83 / 29.48 1 6.24 1 / 1
ActivityNet 17031 / 4885 40.25 / 118.20 1 12.02 1 / 1
TACoS 4083 / 25 31.87 / 367.15 1 8.53 1 / 1
$\text{MVMR}_{\text{Charades-STA}}$ 3716 / 1334 7.83 / 29.48 3.07 6.23 5 / 50
$\text{MVMR}_{\text{ActivityNet}}$ 16941 / 4885 40.32 / 118.20 1.11 12.03 5 / 50
$\text{MVMR}_{\text{TACoS}}$ 2055 / 25 28.55 / 367.15 2.24 7.31 5 / 5
Table 1: Summary of datasets. Remarkably, the queries of $\text{MVMR}_{\text{Charades-STA}}$ and $\text{MVMR}_{\text{TACoS}}$ generally include multiple positive moments, and it proves that existing VCMR studies are exposed to the risk of false negatives.
2.2 Video Corpus Moment Retrieval
Existing studies (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022) have tried to extend the NLVL to search on multiple video pools, naming it the Video Corpus Moment Retrieval (VCMR) task, which focuses on retrieving a moment from multiple videos for a given query. However, the VCMR simply extends the NLVL task by designating only one matched positive video moment for a target query as a positive, while labeling all unmatched videos as negative videos. This approach deviates from the practical scenario of the video moment search since multiple positive videos can be present for a target query. Moreover, existing VCMR studies have failed to consider the possibility of false-negative when constructing the negative video set for a specific query, which poses a significant reliability issue.
3 Problem Definition
MVMR task evaluates video moment retrieval models under an MVMR dataset with $k$ positive videos and $n-k$ negative videos for each query. Given a positive video set $V^{+}_{q}=\{v^{+}_{1},...,v^{+}_{k}\}$ and a negative video set $V^{-}_{q}=\{v^{-}_{1},...,v^{-}_{n-k}\}$, the MVMR task aims to localize a moment $(x_{s},x_{e})$ of a specific video $v$ that matches the query from massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$, where $x_{s}$ and $x_{e}$ mean start and end points of the moment, respectively. In MVMR task, we aim to retrieve an optimal positive moment $(x^{*}_{s},x^{*}_{e})$ that matches semantically with the text query $q$. We derive confidence scores of moments for the whole $n$ massive videos and select the moment with the highest score as a prediction.
4 MVMR Dataset Construction
Our MVMR setting automatically reconstructs an NLVL dataset by expanding it to associate each query with multiple positive and negative moments. The paired positive and negative videos are carefully selected with semantic similarity check techniques. In the following, we detail our process for constructing an MVMR dataset and present the analyses of the three constructed MVMR datasets.
4.1 Semantic Similarity Check
This section examines the similarity between a target query $q$ and a video $v$ using semantic text embedding and video captioning evaluation models.
Query Similarity-based Check.
NLVL datasets consist of video set $V=\{v_{1},...,v_{l}\}$ and paired query set $Q_{v}=\{q_{1},...,q_{m}\}$ for each video $v$, where $l$ and $m$ is the number of videos and text queries included in a dataset, respectively. Since each text query semantically describes the paired moment, we use queries to obtain semantic information about the paired video. For example, suppose that a video $v$ is paired with a query $q$, "The person gets out a knife", in a moment $(x_{s},x_{e})$. Since a target query $\hat{q}$, "The person takes out a knife", is highly similar to the query $q$, $\hat{q}$ is associated with the moment $(x_{s},x_{e})$ of the video $v$. We use a semantic text embedding model, SimCSE Gao et al. (2021), to distinguish positive and negative videos for each query. Specifically, we first obtain embeddings $e_{\hat{q}}$ and $e_{q_{1}},...,e_{q_{m}}$ for the target text query $\hat{q}$ and all video text queries $q_{1},...,q_{m}$ using SimCSE, respectively. We calculate cosine similarity $s_{te}(\hat{q},q_{i})$ between $e_{\hat{q}}$ and $e_{q_{i}}$ to quantify the semantic matching, and select the maximum similarity among $s_{te}(\hat{q},q_{1}),...,s_{te}(\hat{q},q_{m})$ to use it as the query-video matching score $s_{te}(\hat{q},v)$ of the target query $\hat{q}$ and the video $v$. We select max aggregation to obtain $s_{te}(\hat{q},v)$ since the video should be regarded as a positive video, even if only one similar query to a target query exists in a video.
Figure 2: Examples of constructed MVMR datasets. We visualize positive and negative video sets for the query “person closes the door" of the $\text{MVMR}_{\text{Charades-STA}}$. A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive. The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
Query-video Similarity-based Check.
Queries associated with a video may describe the video limitedly since they may not cover all of its context. Therefore, we examine a visual semantic matching between a query and a video to filter positive and negative videos. We use EMScore Shi et al. (2022), a video captioning evaluation model, to quantify the visual semantic similarity $s_{tv}$, used as a complementary filtering method on videos primarily filtered through the query similarity $s_{te}$.
4.2 Distinguishing Positive and Negative
This section describes constructing positive and negative video sets using the calculated query-video similarity, $s_{te}$ and $s_{tv}$. Positive candidates are distinguished by primarily excluding videos $s_{te}(q,v)<t_{te}^{+}$, where $t_{te}^{+}$ is a hyper-parameter. We further exclude videos that have lower similarity than the mean $s_{tv}$ of all golden positive query-video pairs $(q,v^{*})$ for all queries. We construct final MVMR positives by randomly sampling $k$ videos from this filtered set of positive candidates. Negative candidates are defined by excluding videos with $s_{te}(q,v)>t_{te}^{-}$, where $t_{te}^{-}$ is also a hyper-parameter. We additionally exclude videos that have higher similarity than the mean $s_{tv}$ of query-video pairs $(q,v^{-})$, which are primarily filtered as negatives using $s_{te}$. Like MVMR positives, we construct final MVMR negatives by randomly sampling $(n-k)$ videos from this set of negative candidates. From this process, a total of $n$ videos retrieval pool $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ is obtained for each query.
5 MVMR Dataset Analysis
5.1 MVMR Settings.
This section describes the used NLVL datasets and detailed settings to derive MVMR positive and negative candidates using SimCSE and EMScore.
NLVL Datasets.
We extend widely-used NLVL datasets (Charades-STA Gao et al. (2017), ActivityNet Krishna et al. (2017), and TACoS Regneri et al. (2013)) to construct three MVMR evaluation datasets, and name the constructed datasets as $\text{MVMR}_{\textsc{\{\$datasource\}}}$. Each NLVL dataset consists of multiple videos with query-moment pairs. The details of each NLVL dataset are shown in Appendix A.
Filtering Settings.
We use STS16 dataset Agirre et al. (2016) to empirically validate varying SimCSE filtering hyper-parameters, $t^{+}_{te}$ and $t^{-}_{te}$. The dataset provides text pairs and human-annotated similarity scores (five-point Likert scale); two texts are regarded as similar when the score is higher than 3. We analyze the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ to derive MVMR positive candidates since it is regarded as the best threshold to distinguish similar queries. We also select $t^{-}_{te}=0.5$ to filter MVMR negative candidates. The distribution of SimCSE and human-annotated scores are shown in Appendix B.
Figure 3: Qualitative Analysis for Filtering Methods. We visualize the derived similarity scores of SimCSE and EMScore to verify the constructed MVMR datasets. We use T-SNE to reduce the dimension of each query embedding for displaying each query (dot) of SimCSE Similarity Distribution.
5.2 MVMR Analysis
We set a video retrieval pool as $n$, containing at most $k$ positive videos for each query. If the number of MVMR positive candidates derived for each query is less than $k$, we include all positive candidates as positive samples. Also, if the total number of MVMR positive and MVMR negative candidates for a query is less than $n$, the query is excluded from the final MVMR dataset.
Datasets Statistic.
We set the number of videos in the $\text{MVMR}_{\text{Charades-STA}}$ and the $\text{MVMR}_{\text{ActivityNet}}$ retrieval pools as $n=50$, and the maximum number of positive moments per query $k$ is set to $5$. The former consists of 3,716 queries for 1,334 videos, each containing an average of 3.07 positive moments. The latter contains 16,941 queries for 4,885 videos, each including an average of 1.11 positive videos. We set the number of videos for the $\text{MVMR}_{\text{TACoS}}$ and the maximum number of positive moments per query as $5$ and $5$, respectively, since TACoS test set contains only 25 videos. It includes 2,055 queries for 25 videos, each containing an average of 2.24 positive videos. The summary of the MVMR datasets is shown in Table 1.
Dataset Quality.
[…]
Figure 4: Cohesiveness of the Positive and Negative sets. The X and Y axes mean the word co-occurrence ratio and its frequency, respectively. The closer the word’s co-occurrence ratio approaches 1.0, the more cohesive it is.
Human Evaluation.
We recruit crowd workers and ask them to examine 100 queries manually and their paired videos (by our method) and answer whether they are correctly labeled. Overall, we analyze 4,900, 4,900, and 400 videos for Charades-STA, ActivityNet, and TACoS, respectively, since each MVMR dataset consists of 49, 49, and 4 newly added videos (either positive or negative samples). The original video (i.e., golden positive sample) is excluded from this investigation. From human evaluation, we confirm that the newly introduced dataset contains only 1.5%, 0.2%, and 3.5% falsely categorized videos for the three datasets, respectively. TACoS exhibits a relatively higher query-relevant video ratio than the other two datasets since its test set includes only 25 videos and consists of considerably similar cooking activities within the kitchen.
Analysis on Semantic Filtering Methods.
We visualize the scores derived using SimCSE and EMScore to analyze our MVMR datasets qualitatively. Figure 3-(a) displays SimCSE embeddings of all queries in each dataset, illustrating the similarity between all queries and a specific target query, "person turn a light on." (left) and "He is using a push mower to mow the grass." (right). The figure of TACoS can also be found in Appendix B. The figures show that queries exhibit well-defined clusters for Charades-STA and TACoS, revealing the effectiveness of our query similarity-based filtering. These results are also supported by Figure 3-(b) SimCSE score histogram, showing that Charades-STA and TACoS have many similar queries. Figure 3-(c) illustrates histograms of EMScore similarity between queries and videos. The red and blue histograms correspond to the EMScore distribution of positive and negative samples, respectively. For the ActivityNet, EMScore effectively distinguishes positives and negatives. This success is attributed to that ActivityNet includes detailed queries and videos with diverse features. According to the findings presented in this section, the filtering method’s efficacy varies depending on the dataset’s characteristics. Consequently, a combination of two filtering methods (i.g., SimCSE and EMScore) should be employed in a complementary manner to ensure the construction of a reliable MVMR dataset.
6 Methods: CroCs
Figure 5: CroCs Overview. We adopt informative negative contrastive learning to solve the MVMR task. The dots and triangles are the features of moments and texts. The blue dash lines are matched moment-text pairs to be pulled in, while the red dash lines are negative samples of intra/inter-video to be pushed away. The yellow and orange dash lines are unmatched moment-text pairs, but not to train by filtering out since they are easy and false-negatives.
This section introduces a novel contrastive learning method, CroCs, which stands for Cross-directional Hard and Reliable Negative Contrastive Learning, to solve the MVMR task. A key challenge in solving the MVMR task is distinguishing positives from multiple negative distractors. Therefore, we select MMN Wang et al. (2022) since it is the optimal baseline model for the MVMR task as it uses contrastive learning to distinguish negative samples from positives effectively. We use the MMN as the backbone of our model and enhance its ability by applying CroCs in training procedures to solve the challenge in the MVMR task. MMN has limitations in neglecting the potential presence of false-negative video moments since it regards all randomly selected in-batch samples as negatives in contrastive learning. Furthermore, it overlooks the importance gap among negative samples. Our CroCs solves these two limitations of the MMN by adopting two contrastive learning procedures: (1) Weakly-supervised potential negative learning and (2) Informative hard-negative learning.
6.1 Weakly-supervised Contrastive Learning
A critical consideration in the contrastive learning framework is the potential presence of false-negatives. This issue arises when all videos, except for the golden video, are categorized as negative training samples. To mitigate this concern, we propose excluding false-negative videos to refine the contrastive learning process. Specifically, we adopt the SimCSE filtering to identify and exclude false-negative videos as follows:
$V^{-}_{tr}(q)=\{v|s_{te}(q,v)<t_{tr}\},\>\>\ \forall{v}\in V\>$ (1)
where $V^{-}_{tr}(q)$ is a negative sample set of a specific query $q$ to use during the contrastive learning procedure. $s_{te}(q,v)$ is query-video similarity described in Section 4.1. $t_{tr}$ is a hyper-parameter for SimCSE score filtering. If a specific video shows a value of $s_{te}(q,v)$ above the threshold $t_{tr}$ for target query $q$, we classify $v$ as false-negative for $q$. Therefore, we exclude $v$ in negative samples of contrastive learning for $q$.
6.2 Hard-negatives Contrastive Learning
Easy or false-negatives may distract a model from learning proper knowledge during the training process. Inspired by recent progress in using hard-negative sampling for the document retrieval task Zhou et al. (2022), we categorize negatives according to the matching score between a target query and negative moments as follows: (1) Negatives that are clearly irrelevant and have low matching scores should be sampled less frequently (Easy-negative); (2) Negatives that are highly relevant and have high matching scores should also be sampled less frequently (False-negative); (3) Negatives that are uncertain and have matching scores similar to true-positives should be sampled more frequently since they provide useful information (Hard-negative). Based on these criteria, we distinguish hard-negative samples and adopt a retraining process for a model already trained on a specific dataset using only hard-negative samples. We define the negative moments and query sampling distribution as follows:
$p^{\prime}(m|q)\propto\exp{(-(r(q,m)-\overline{r}(q,m^{+}))^{2})},\>\ \forall m\in M^{-}_{q}$ (2)
$p^{\prime}(q|m)\propto\exp{(-(r(q,m)-\overline{r}(q^{+},m))^{2})},\>\ \forall q\in Q^{-}_{m}\>\>$
where $r(q,m)$ is a matching score between a query and a moment calculated using the already trained video moment retrieval model, and $\overline{r}(q,m^{+})$ and $\overline{r}(q^{+},m)$ is mean matching scores for all golden positive moments and queries, respectively. $M^{-}_{q}$ and $Q^{-}_{m}$ is potential negative candidates for a query $q$ and a moment $m$, respectively. $p\textquoteright(m|q)$ means the probability that moment $m$ is sampled as negative when given query $q$. Therefore, negative candidate moments with a close distance to the average matching score of golden query-moment pairs have a high sampling probability. Likewise, $p\textquoteright(q|m)$ means the probability that query $q$ is sampled as negative when given moment $m$. We also reformulate $p\textquoteright(m|q)$ to $p\textquoteright(v|q)$ for computational efficiency in a sampling procedure as follows:
$r(q,v)=\max{(\{r(q,m_{1}),...,r(q,m_{l})\})},\>\>\ \forall m_{i}\in v\>\>\>\>\>\>$ (3)
$p^{\prime}(v|q)\propto\exp{(-(r(q,v)-\overline{r}(q,v^{+}))^{2})},\>\>\ \forall v\in V^{-}_{q}\>\>\>\>\>\>$
$p^{\prime}(q|v)$ can be formulated similarly by calculating $\overline{r}$ about positive queries for each video.
6.3 Reliable Mutual Matching Network
This section describes a model architecture to apply the proposed contrastive learning method, CroCs. We adopt a bi-encoder architecture with a late modality fusion by an inner product in the joint visual-text representational space following the baseline MMN. By adopting a bi-encoder architecture, we can efficiently search for moments with pre-calculated video moment representations given a user query. Our whole model architecture is shown in Figure 5. Our model represents each query and video moment using query and video encoders to two joint visual-text spaces (IoU and mutual matching spaces). From the joint visual-text space, we compute the matching score between the query and video moment representations and select the best video moment. The details of the model architecture are described in the appendix C.1.
We select the binary cross entropy and contrastive learning losses and train our model following two steps: (1) the first training step using the two loss functions (binary cross entropy and contrastive learning losses) filtering out false-negatives described in Section 6.1; (2) the second training step using two loss functions with only hard-negatives described in Section 6.2. To calculate the similarity between the text and video moment features projected in the IoU space, we compute the cosine similarity $s^{iou}$. To adjust the range of the final prediction, we multiply the cosine similarity by a factor of 10, following the MMN. This amplification results in the final prediction $p^{iou}_{i}=\sigma(10\cdot s^{iou}_{i})$, where $\sigma$ represents the sigmoid function. The binary cross-entropy loss is then calculated as follows:
$L_{bce}=-\frac{1}{C}\sum^{C}_{i=1}{(y_{i}\log{p^{iou}_{i}}+(1-y_{i})\log{(1-p^{iou}_{i})})}$ (4)
where $p^{iou}_{i}$ is the confidence score of each moment, and $C$ is the total number of valid candidates. We calculate $L_{bce}$ for positive and true-negative samples derived in Section 6.1.
We use the cross-directional contrastive learning loss to effectively train the model using the informative negatives derived from the methods described in Section 6.1 and 6.2. The mutual matching contrastive learning loss considering the informative negative samples selection process is as follows:
$p^{*}(q_{i}|m)=\frac{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})}{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})+\sum_{j\in\mathcal{I}_{m}^{-}}{\exp(f_{j}^{q\mathsf{T}}f^{m}/\tau_{m})}}\>\>\>$ (5)
$p^{*}(m_{i}|q)=\frac{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})}{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})+\sum_{j\in\mathcal{I}_{q}^{-}}{\exp(f_{j}^{m\mathsf{T}}f^{q}/\tau_{q})}}\>\>\>\>\>$
$L_{rmm}=-(\sum^{N}_{i=1}{(\log{p^{*}(i_{m}|q_{i})}})+\sum^{N}_{i=1}{(\log{p^{*}(i_{q}|m_{i})}}))\>\>\>\>\>\>\>\>\>\>\>$
Dataset Model R@1 R@1 R@1 R@5 R@5 R@5 R@20 R@20 R@20 R@50 R@50 R@50
IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7
$\text{MVMR}_{\text{Charades-STA}}$ CET 3.96 2.96 2.02 15.34 11.79 6.97 39.85 32.21 19.91 59.58 50.54 33.88
BET 4.57 2.99 1.94 14.26 10.98 7.00 31.05 25.11 16.55 45.88 37.86 26.08
G2L 4.30 2.78 1.97 14.34 10.59 7.36 40.89 34.74 21.33 71.92 62.72 44.82
MMN 13.99 12.78 8.40 37.86 33.29 22.36 64.80 57.83 41.71 80.11 73.09 56.75
${}^{\dagger}$ 18.38 16.31 10.76 43.86 39.05 26.99 69.48 62.89 46.39 82.88 76.08 60.28
CroCs 19.27 17.38 11.44 43.54 39.02 27.91 68.08 60.98 46.39 82.99 76.29 60.76
$\text{MVMR}_{\text{ActivityNet}}$ CET 1.26 0.90 0.56 2.81 2.32 1.75 5.41 4.73 3.87 8.97 8.02 6.56
BET 1.56 1.16 0.76 3.02 2.55 2.03 5.30 4.64 3.80 8.13 7.10 5.88
G2L 1.66 1.32 0.96 7.82 6.43 4.55 26.84 21.66 14.81 62.99 50.07 32.99
MMN 13.82 10.71 6.93 31.54 25.04 16.55 54.79 44.50 30.62 73.08 62.09 45.76
${}^{\dagger}$ 16.46 12.68 7.87 36.03 28.47 18.16 57.91 46.52 31.12 73.82 62.12 44.89
CroCs 20.63 15.58 9.51 44.13 34.70 22.40 67.01 55.58 38.55 79.91 70.02 53.20
$\text{MVMR}_{\text{TACoS}}$ CET 8.95 5.79 3.89 26.18 18.88 12.17 54.84 42.53 23.70 - - -
BET 8.22 5.30 2.97 29.46 14.84 8.32 55.67 39.76 21.51 - - -
G2L 8.72 5.82 4.31 30.61 21.51 12.29 56.52 42.35 22.52 - - -
MMN 10.56 8.61 5.60 36.16 25.84 15.47 58.20 43.89 23.80 - - -
${}^{\dagger}$ 13.43 10.75 6.76 36.45 27.01 15.28 58.30 44.38 25.16 - - -
CroCs 13.24 10.75 7.30 39.22 29.68 16.50 60.05 44.77 24.62 - - -
Table 2: Experimental Results on MVMR datasets. Bolded and under-lined results indicate the $1^{st}$ and $2^{nd}$ best performance, respectively. ${}^{\dagger}$ is the model using only weakly-supervised contrastive learning (§6.1). CroCs uses both of two contrastive learning methods (§6.1 and §6.2). We train all the baselines as three trials and report the averaged MVMR scores.
where $q_{i}$ and $m_{i}$ are corresponding positive query and moment for each moment and query, respectively. $f^{m}$ and $f^{q}$ are the moment and query features in the joint visual-text space. $\tau_{q}$ and $\tau_{m}$ are temperatures. $\mathcal{I}^{-}$ means the indices of informative negatives derived using our sampling methods.
The final loss, denoted as $L$, is formulated as a linear combination of the binary cross-entropy loss and the reliable mutual matching loss. The matching score $s$, for a specific moment given the text query, is obtained by multiplying the iou score $s^{iou}$ with the reliable mutual matching score $s^{rmm}$. The reliable mutual matching score is calculated for the features in the mutual matching space in the same way as the iou score.
$L=L_{bce}+\lambda L_{rmm},\>\>\>\>\ s=s^{iou}\cdot s^{rmm}$ (6)
7 Experiments
7.1 Experimental Settings
This section describes the experimental settings of the CroCs and other baselines. More details are shown in Appendix C.
Implementation Details.
We use standard off-the-shelf video feature extractors without any fine-tuning: VGG Simonyan and Zisserman (2015) feature for Charades-STA; C3D Tran et al. (2015) feature for ActivityNet and TACoS following Wang et al. (2022). We set the filtering hyper-parameter as $t_{tr}=0.9$ for the weakly-supervised contrastive learning. We choose the number of the CroCs hard negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. These hyper-parameters are chosen from varying parameter setting experiments as shown in Figure 8. In the hard-negative fine-tuning stage, the learning rates are degraded by multiplying 0.1.
Baselines.
We first select the state-of-the-art models, MMN Wang et al. (2022) and G2L Li et al. (2023), as strong baselines. As the MVMR  task can be formulated as an open-domain question answering (QA) task, we further choose two types of QA models that employ a cross-encoder architecture (i.e., Cross-Encoder Transformer) and a dual-encoder architecture (i.e., Bi-Encoder Transformer), respectively. These architectures are widely adopted in existing NLVL tasks Zhang et al. (2020b); Gao and Xu (2021) due to the closeness of the two tasks.
Evaluation Metrics.
We evaluate baselines by calculating Rank n@m. It is defined as the percentage of queries having at least one correctly retrieved moment, i.e., IoU $\geq m$, in the top-$n$ derived moments. The existing NLVL studies have reported the results of $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5\}$ Wang et al. (2022). But, since our MVMR setting searches for massive videos, we report the extended results as $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20,50\}$ for Charades-STA and AcitivityNet and $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20\}$ for TACoS.
7.2 MVMR Results
MVMR Performance.
We evaluate the performance of CroCs and the baseline models on the introduced MVMR datasets. Table 2 shows the experimental results of CroCs and other baselines for our MVMR task. Surprisingly, although CET and BET show quite strong performance on the NLVL task, their performance on the MVMR task notably drops. For examples, CET and MMN show average 34.5 and 29.7 degradation on R@1(IOU0.5) for all MVMR datasets. These outcomes indicate that the current NLVL task performance alone is not enough to verify that a model is well-trained. We reveal that CroCs outperforms all other baselines on MVMR task, and this suggests that our model correctly discriminates features of negative and positive moments. Experiments show that the models on Charades-STA and TACoS significantly improve performance from the weakly-supervised learning since similar queries exist frequently in these datasets and the possibility of false-negatives is high. However, the performance of ActivityNet is attributed to the hard-negative learning since it includes various and complex text queries.
MVMR Settings with Video Retrieval.
The practical usage scenario for the MVMR should include video retrieval. Therefore, We also conduct MVMR experiments with a video retrieval model, PRMR model Dong et al. (2022), shown as Table 3. We construct a pipeline to filter top-5 logit videos from our MVMR retrieval pool using the PRMR model and solve MVMR task. We use the publicly deployed PRMR model on Charades-STA and ActivityNet datasets. These experiments prove that video retrieval models are helpful to increase MVMR performance, and CroCs still shows superior performance than MMN in this pipeline.
Model R@1 IOU0.3 R@1 IOU0.5 R@5 IOU0.3 R@5 IOU0.5
$\text{MMN}_{\text{Charades-STA}}$ 32.15 27.10 65.97 57.66
$\text{CroCs}_{\text{Charades-STA}}$ 36.64 29.27 68.06 59.78
$\text{MMN}_{\text{ActivityNet}}$ 32.39 25.04 66.84 55.03
$\text{CroCs}_{\text{ActivityNet}}$ 39.25 29.59 71.54 60.02
Table 3: MVMR Experiments with Video Retrieval.
8 Conclusion
In this paper, we propose the Massive Videos Moment Retrieval (MVMR) task, which aims to detect a moment for a natural language query from a massive video set. To stimulate the research, we propose a simple automatic method to construct reliable MVMR datasets and build three MVMR datasets using existing NLVL datasets. We introduce a robust training method called CroCs to distinguish positives from negative distractors effectively in solving MVMR task.
Limitations
Our dataset construction framework automatically defines positive and negative sets using semantic similarity evaluation. Unlike the existing VCMR framework, which overlooks the risk of labeling errors in its automated dataset construction process, our MVMR framework introduces some filtering methods to mitigate labeling errors. However, our method still faces limitations due to its reliance on automatic labeling, which fundamentally carries a risk of labeling inaccuracies. To eliminate the risk of labeling errors thoroughly, a constructed dataset should be reviewed manually through complete enumeration.
Appendix A Datasets Details
Charades-STA
is an extended dataset of action recognition and localization dataset Charades Sigurdsson et al. (2016) by Gao et al. (2017) for video moment retrieval. It is comprised of 5,338 videos and 12,408 query-moment pairs in the training set, and 1,334 videos and 3,720 query-moment pairs in the test set. We report evaluation results on test set in our experiments.
ActivityNet-Captions
is constructed on ActivityNet v1.3 dataset Heilbron et al. (2015), where included videos cover various complex human actions. It is originally designed for video captioning, and recently used into video moment retrieval. It contains 37,417, 17,505, and 17,031 query-moment pairs for training, validation, and testing respectively. We report the evaluation result on val_2 set following the setting of the MMN paper Wang et al. (2022).
TACoS
includes 127 videos selected from the MPII-Cooking dataset. It consists of 18,818 query-moment pairs of various cooking activities in the kitchen annotated by Regneri et al. (2013). It Gao et al. (2017) consists of 10,146, 4,589, and 4,083 query-moment pairs for training, validation and testing, respectively. We report evaluation results on test set in our experiments.
Appendix B Analysis on Filter Methods
We construct reliable MVMR datasets using SimCSE and EMScore. The visualizations of the similarity scores derived by SimCSE and EMScore is shown as Figure 7. For the SimCSE distribution, we select target queries of "person turn a light on", "The person peels a kiwi", and "He is using a push mower to mow the grass" for the Charades-STA, TACoS, and ActivityNet, respectively.
Figure 6 shows the distribution of SimCSE and human-annotated STS16Agirre et al. (2016) dataset scores. STS16 dataset provides text pairs and annotated similarity scores (five-point Likert scale). In the dataset, two texts are regarded as similar when the score is higher than 3. We visualize the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ and $t^{-}_{te}=0.5$ as thresholds for the SimCSE filtering to derive MVMR positive and negative candidates, respectively. STS16 dataset includes text pairs that are challenging to distinguish. However, the three datasets utilized in our experiment do not demand similarity measurement at such a complex level. Consequently, we can reliably discern the similarity between queries using the provided thresholds.
Figure 6: The distribution of SimCSE and Human-annotated Scores. The X and Y axes mean the human-annotated and the SimCSE scores, respectively. The width of each graph corresponds to the number of samples.
[…]
Figure 7: The Visulization of the SimCSE and EMScore Similarity Distribution.
Appendix C Implementation Details
We use standard off-the-shelf video feature extractors without any fine-tuning. We use VGG Simonyan and Zisserman (2015) feature for Charades-STA and C3D Tran et al. (2015) feature for ActivityNet and TACoS following the previous work Wang et al. (2022). We train CroCs and MMN on 1 NVIDIA RTX A6000 GPU and early stop by averaging scores of all evaluation metrics.
C.1 CroCs
We implement CroCs by following MMN’s model architecture. CroCs represents each query and video moment using query and video encoders to a joint visual-text space. From the joint visual-text space, CroCs computes the similarity between the query and video moment representations and selects the best-matched video moment.
Query and Video Encoders.
We choose DistilBERT (Sanh et al., 2019) for the text query encoder following MMN since it is a lightweight model showing significant performance. We calculate the representation of each query $f^{q}\in\mathbb{R}^{d}$ using the global average pooling over all tokens. We adopt the approach of encoding the input video as a 2D temporal moment feature map, inspired by 2D-TAN Zhang et al. (2020b) and MMN. To achieve this, we segment the input video into video clips denoted as $\{c_{i}\}_{i=1}^{l_{c}/k}$, where each clip $c_{i}$ consists of $k$ frames. $l_{c}$ means the total number of frames in the video. The clip-level representations are extracted using a pre-trained visual model (e.g., C3D). By sampling a fixed length with a stride of $\frac{l_{c}}{k\cdot N}$, we obtain $N$ clip-level features. These features are then passed through an FC layer to reduce the dimensionality, resulting in the features $\{f^{v}_{i}\}^{N}_{i=1}$, where $f^{v}_{i}\in\mathbb{R}^{d}$. Utilizing these features, we construct a 2D temporal moment feature map $F\in\mathbb{R}^{N\times N\times d}$ by employing max-pooling as the moment-level feature aggregation method following the baseline (Wang et al., 2022). Additionally, we generate a 2D feature map $F^{\prime}$ of the same size by passing $F$ to 2D Convolution, allowing the representation of moment relations as employed in MMN.
Joint Visual-Text Space.
Initially, we apply layer normalization to both the video moments and text features. Subsequently, we utilize a linear projection layer and a 1$\times$1 convolution layer to project the text and video in the same embedding space, respectively. The projected features are then employed in two distinct representational spaces: the IoU space and the mutual matching space. These spaces serve as the basis for computing the binary cross-entropy loss and the contrastive learning loss, respectively.
Hyperparameter Settings.
Our convolution network for deriving 2D features is exactly the same as MMN, including the number of sampled clips $N$, the number of 2D conv layers $L$, kernel size, and channels. We set the dimension of the joint feature visual-text space $d=256$, and temperatures $\tau_{m}=\tau_{q}=0.1$. We utilize the pre-trained HuggingFace implementation of DistilBERT Wolf et al. (2019). We set margin $\psi$ as 0.4, 0.3, and 0.1 for Charades-STA, ActivityNet, and TACoS, respectively. We use AdamW Loshchilov and Hutter (2017) optimizer with learning rate of $1\times 10^{-4}$, $8\times 10^{-4}$, $1.5\times 10^{-3}$ for Charades-STA, ActivityNet, and TACoS, respectively. We set $\lambda$ as 0.05 for Charades-STA and TACoS and 0.1 for ActivityNet. We set the reliable true-negatives filtering threshold $t_{tr}=0.9$. We set the number of hard-negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively. The experimental results for various hyper-parameters for the hard-negative sampling are shown in Figure 8.
Figure 8: Evaluation results for various hyper-parameters. We search various hyper-parameters for the hard negative sampling. We measure R@1 (IoU>m), where m$\in\{0.3,0.5,0.7\}$ to find the best hyper-parameters. In this figure, each $(x,y)$ means the number of negative samples. $x$ and $y$ mean a number of negative samples for $p^{\prime}(v|q)$ (query-to-video) and $p^{\prime}(q|v)$ (query-to-video), respectively. Consequently, we select the number of the CM-ANS samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
C.2 Baselines
MMN
Wang et al. (2022) adopts a dual-modal encoding design to get video clip and text representations. We utilized the publicly deployed code [^2] https://github.com/MCG-NJU/MMN to implement our own MMN model for our experiment using exactly the same hyper-parameters setting.
G2L
Li et al. (2023) uses a semantically aligned and uniform video grounding framework via geodesic and game theory. We utilized the publicly deployed code [^3] https://github.com/lihxxxxx/G2L?tab=readme-ov-file to implement our own G2L model for our experiment using exactly the same hyper-parameters setting proposed by the paper.
Cross-Encoder Transformer (CET)
is a cross-modal encoding model implemented based on a 6-layered transformer architecture. It is designed to predict the start and end points of a video moment by utilizing a model architecture, which is used in the QA task. We concatenate the pre-extracted video clip features and the text features derived from the Distil-BERT to use as an input of the transformer. To concatenate two extracted features, we should equalize the dimensions of two features with a linear layer as follows:
$\bar{H}^{v}=H^{v}W^{v}$ (7)
$\bar{H}^{q}=H^{q}W^{q}$
where $H^{v}\in\mathbb{R}^{l_{v},d^{v}}$, $H^{q}\in\mathbb{R}^{l_{q},d^{q}}$, $W^{v}\in\mathbb{R}^{d^{v},d}$, $W^{q}\in\mathbb{R}^{d^{q},d}$. $H^{v}$ is the pre-extracted video clip features and $H^{q}$ is the features extracted using Distil-BERT. And then, we derive the representation of each video clip feature from the last layer of the transformer-encoder as follows:
$H^{v,q}=TransformerEnc(\bar{H}^{v}\oplus\bar{H}^{q})$ (8)
where $H^{v,q}\in\mathbb{R}^{l_{vq},d}$, $l_{vq}=l_{v}+l_{q}$. We utilize only the representation part of video features, and the 3-layered MLPs and Sigmoid follow them to predict the start and end positions of the moment, respectively, as follows:
$p^{s}=Sigmoid(MLP_{s}(H^{v,q}_{1:l_{v}}))$ (9)
$p^{e}=Sigmoid(MLP_{e}(H^{v,q}_{1:l_{v}}))$
where $p^{s}$ and $p^{e}\in\mathbb{R}^{l_{v}}$. We also introduce the interpolation labels $y_{i}^{s}$, $y_{i}^{e}$ to utilize abundant information of moment labels as follows:
$y_{i}^{s}=\left\{\begin{array}[]{ c l }({\frac{x_{e}-i}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>\>$ (10)
$y_{i}^{e}=\left\{\begin{array}[]{ c l }({\frac{i-x_{s}}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>$
We finally use the binary cross-entropy to calculate a loss for the pair of $(p_{i}^{s},y_{i}^{s})$ and $(p_{i}^{e},y_{i}^{e})$ as follows:
$\mathcal{L}^{s}=BCE(p^{s},y^{s})$ (11)
$\mathcal{L}^{e}=BCE(p^{e},y^{e})$
where BCE means the binary cross-entropy. The final loss to train the CET model is defined as follows:
$\mathcal{L}=\frac{\mathcal{L}^{s}+\mathcal{L}^{e}}{2}$ (12)
We use $k=10$ and $d=512$ as a hyperparameter and AdamW optimizer with learning rate 1e-3.
Bi-Encoder Transformer (BET)
is a dual-modal encoding model that encodes text and video clip features independently. The pre-extracted video clip features are encoded using a 6-layered transformer, and text features are encoded using the Distil-BERT, followed by the average-pooling function and 3-layered MLPs.
$\tilde{H}^{v}=TransformerEnc(\bar{H}^{v})$ (13)
$\tilde{H}^{s}=MLP_{s}(Pool(\bar{H}^{q}))$ (14)
$\tilde{H}^{e}=MLP_{e}(Pool(\bar{H}^{q}))$
Where $\tilde{H}^{v}\in\mathbb{R}^{d}$, $\tilde{H}^{s}\in\mathbb{R}^{d}$ and $\tilde{H}^{e}\in\mathbb{R}^{d}$. $Pool$ means the average pooling. We calculate the cosine similarity between the encoded video clip representations and the encoded start and end representations to predict a moment.
$p^{s}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{s}))$ (15)
$p^{e}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{e}))$
We use the interpolation labels and the binary cross-entropy loss for the BET, similar to CET. We use $r=10$ and $d=512$ as a hyperparameter and AdamW optimizer with learning rate 1e-3.
Appendix D Examples of False Negative Videos
Figure 9: Examples of constructed MVMR datasets. We visualize positive and negative video sets for the query “person closes the door" of three constructed datasets. A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive. The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
Illustrative instances of constructed MVMR datasets can be found in Figure 9. This figure shows video instances that have been identified as similar to a specific target query. The target queries correspond to queries of Charades-STA, ActivityNet, and TACoS datasets, sequentially.
Appendix E Human Evaluation for the constructed MVMR Datasets
Figure 10: Human Evaluation Sheet Example.
We conduct a human evaluation on our constructed MVMR datasets to reveal that our method construct MVMR datasets effectively. We recruit crowd workers fluent in English through the university’s online community. The recruited annotators were provided with a detailed description of task definitions, instructions. An example of a sheet we use to conduct a human evaluation is shown in Figure 10.

### v4

@@
-Is it Really Negative? Evaluating Natural Language Video Localization Performance on Multiple Reliable Videos Pool
-Abstract
-With the explosion of multimedia content in recent years, Video Corpus Moment Retrieval (VCMR), which aims to detect a video moment that matches a given natural language query from multiple videos, has become a critical problem.
-However, existing VCMR studies have a significant limitation since they have regarded all videos not paired with a specific query as negative, neglecting the possibility of including false negatives when constructing the negative video set.
-In this paper, we propose an MVMR (Massive Videos Moment Retrieval) task that aims to localize video frames within a massive video set, mitigating the possibility of falsely distinguishing positive and negative videos.
-For this task, we suggest an automatic dataset construction framework by employing textual and visual semantic matching evaluation methods on the existing video moment search datasets and introduce three MVMR datasets.
-To solve MVMR task, we further propose a strong method, CroCs, which employs cross-directional contrastive learning that selectively identifies the reliable and informative negatives, enhancing the robustness of a model on MVMR task.
-Experimental results on the introduced datasets reveal that existing video moment search models are easily distracted by negative video frames, whereas our model shows significant performance.
-[^1] We have opened an anonymous GitHub page to make our code and dataset publicly available: link
-1 Introduction
-Enabled by the increased accessibility of video-sharing platforms along with the advancement of networking and storage technology, a vast number of new content is available on the web daily.
-With this huge number of contents, Natural Language Video Localization (NLVL) task, searching for a video moment suitable for a given query, has emerged as an essential problem (Anne Hendricks et al., 2017; Gao et al., 2017).
-However, this line of research narrows down the task to an impractical setting by targeting the search over a single video; thus, it bears limitations in retrieving relevant information over whole media pools.
-Figure 1: MVMR vs.
-VCMR.
-In existing VCMR studies, only a single golden positive video moment is identified as positive, while entire videos not paired with a specific query are regarded as negative.
-Our proposed MVMR filters positive video moments for a query from the whole video set; thus, it covers more practical and generalized use cases with reliable settings.
+MVMR: A New Framework for Evaluating Faithfulness of Video Moment Retrieval against Multiple Distractors
+Abstract.
+With the explosion of multimedia content, video moment retrieval (VMR), which aims to detect a video moment that matches a given text query from a video, has been studied intensively as a critical problem.
+However, the existing VMR framework evaluates video moment retrieval performance, assuming that a video is given, which may not reveal whether the models exhibit overconfidence in the falsely given video.
+In this paper, we propose the MVMR (Massive Videos Moment Retrieval for Faithfulness Evaluation) task that aims to retrieve video moments within a massive video set, including multiple distractors, to evaluate the faithfulness of VMR models.
+For this task, we suggest an automated massive video pool construction framework to categorize negative (distractors) and positive (false-negative) video sets using textual and visual semantic distance verification methods.
+We extend existing VMR datasets using these methods and newly construct three practical MVMR datasets.
+To solve the task, we further propose a strong informative sample-weighted learning method, CroCs, which employs two contrastive learning mechanisms: (1) weakly-supervised potential negative learning and (2) cross-directional hard-negative learning.
+Experimental results on the MVMR datasets reveal that existing VMR models are easily distracted by the misinformation (distractors), whereas our model shows significantly robust performance, demonstrating that CroCs is essential to distinguishing positive moments against distractors.
+Our code and datasets are publicly available: link.
+Figure 1.
+Faithfulness Issue in Video Moment Retrieval.
+If a VMR model is trained without considering overconfidence in a given context (video), it may undergo a Type 1 error since it prioritizes selecting a video moment from the falsely given video (distractor).
+Video Moment Retrieval, Trustworthy AI, Contrastive Learning
+submissionid: 2163journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679838ccs: Information systems Multimedia and multimodal retrieval
+1. Introduction
+Enabled by the increased accessibility of video-sharing platforms along with the advancement of networking and storage technology, a vast amount of new content becomes available on the web daily.
+With this huge number of contents, Video Moment Retrieval (VMR) task, searching for a video moment suitable for a given query, has emerged as an essential problem (Anne Hendricks et al., 2017; Gao et al., 2017).
+However, this line of research evaluates retrieval models assuming that a context (video) is given, it is limited in revealing the capability of distinguishing misinformation.
+Figure 1 exemplifies the problem of faithfulness for misinformation demonstrated by a VMR model.
+This overconfidence issue has been studied recently in the natural language community as a challenging task (Chen et al., 2022; Shi et al., 2023), but it has not been studied sufficiently in the VMR field.
+This problem also arises in a practical video moment retrieval scenario, Video Corpus Moment Retrieval (VCMR), which aims to search for a relevant video moment within a pool of multiple videos (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022).
+The VCMR framework evaluates video moment retrieval performance using a sequential pipeline that integrates video retrieval with video moment retrieval.
+Therefore, it also diverges from evaluating the overconfidence of a VMR model, as the evaluated retrieval performance may predominantly depend on video retrieval capability.
+Figure 2.
+Massive Video Pool Construction.
+From our experiment on the TACoS dataset, if a massive video set is constructed using random sampling, then over 40% of queries include at least one false-negative video, showing the risk of random sampling.
+Our method constructs a massive video set by filtering trustful positive and negative sets using semantic distance verification methods, considering the possibility of false-negatives.
 $v^{+}_{i}$ and $v^{-}_{j}$ mean a positive and a negative video, respectively.
-Therefore, Video Corpus Moment Retrieval (VCMR) task that retrieves a correct positive moment among numerous negative videos has emerged (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022).
-Nonetheless, VCMR also has significant limitations as it merely extends NLVL datasets by categorizing all unpaired moments for a specific text query as negative.
-This assumption diverges from the practical setting of video moment search where multiple positive videos exist in whole media collections.
-Furthermore, existing VCMR studies have overlooked the potential presence of false-negative videos within the unpaired video set when constructing video retrieval pools for a specific query.
-As a result, evaluating the performance of NLVL models using an existing VCMR framework may lead to unreliable results.
-For example, in TACoS test dataset, it is noteworthy that the number of videos with queries exactly matching the text "The person gets out a knife." amounts to 60% of all the videos.
-It is a significant problem since more videos will likely be incorrectly classified as negatives when expanding the matching criteria to include broader semantic similarities.
-To bridge these gaps, we propose an MVMR (Massive Videos Moment Retrieval) task, which expands the search coverage to a massive video set that can include any number of positive videos as shown in Figure 1.
-Specifically, given a query $q$ with a positive videos $v^{+}_{1},...,v^{+}_{k}$ and a negative videos $v^{-}_{1},...,v^{-}_{n-k}$, the MVMR task aims to detect temporal moments in the positive videos that matches the query from a massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$.
-Furthermore, we propose simple and effective methods to construct a practical MVMR dataset by leveraging publicly available NLVL datasets.
-A critical challenge in converting an NLVL dataset into an MVMR dataset is accurately defining positive and negative videos for each query.
-However, manually performing this task by humans is exceedingly labor-intensive due to the massive volume of videos that need to be evaluated for each query.
-To address this issue, we propose an automated framework for MVMR dataset construction.
-This framework employs textual and visual semantic matching evaluation methods to define MVMR positive and negative videos while mitigating the danger of false categorization.
-Specifically, we examine the similarity between a target query and all videos in an NLVL dataset using SimCSE Gao et al.
-(2021) and EMScore Shi et al.
-(2022) models to define a massive video set.
-To this end, we construct three MVMR datasets, and we confirm that three samples of the introduced datasets contain only 1.5%, 0.2%, and 3.5% falsely defined videos, showing that our approach successfully creates a practical benchmark dataset from human evaluation.
-To tackle MVMR task, we also introduce a novel negative sample training method, Cross-directional Hard and Reliable Negative Contrastive Learning (CroCs).
-Specifically, CroCs adopts two training mechanisms: (1) weakly-supervised potential negative learning by assessing the similarity between queries, and (2) hard-negative learning by utilizing the relevance score between queries and moments predicted by a model, which helps find challenging negative samples.
-We demonstrate the performance of strong baseline models, including a state-of-the-art model, to compare with our model in the new scenario (MVMR).
-The experimental results show that the performance of baseline models significantly degrades on the new task.
-For example, the state-of-the-art model, MMN, shows average 29.7% degradation on R@1(IOU0.5) for all MVMR datasets, revealing the difficulty of the MVMR task.
-In contrast, our CroCs outperforms baseline models significantly in the MVMR setting, and this result proves that informative negative samples enhance the performance of the MVMR task.
-Moreover, we solve the MVMR task in a more practical scenario by adopting a pipeline with a video retrieval model, revealing that CroCs also outperforms the strong baseline.
-2 Backgrounds
-2.1 Natural Language Video Localization
-Natural language video localization (NLVL), which aims to detect a specific moment in a video that matches a given text query, is one of the most challenging problems in the video-language multi-modal domain (Zhang et al., 2020b, a; Nan et al., 2021; Gao and Xu, 2021; Liu et al., 2021; Wang et al., 2022).
-Therefore, the NLVL task has been explored extensively, with numerous effective methods emerging.
-Some novel approaches have regarded the problem as question-answering (QA) tasks (Seo et al., 2018; Joshi et al., 2020), by leveraging a QA model to encode a multi-modal representation and then predicting the frames that correspond to the start and end of the moment that match a query (Ghosh et al., 2019; Zhang et al., 2020a).
-However, NLVL studies operate within an impractical framework, focusing on searching within a single video.
-Consequently, they have limitations in retrieving relevant information across the whole media repositories.
+To tackle this issue, we propose an MVMR (Massive Videos Moment Retrieval for Faithfulness Evaluation) task, which expands the search coverage to a massive video set that can include any number of positive and negative videos for evaluating the faithfulness of VMR models.
+Specifically, given a text query $q$ with positive videos $v^{+}_{1},...,v^{+}_{k}$ and negative videos $v^{-}_{1},...,v^{-}_{n-k}$, the MVMR task aims to retrieve temporal moments in the positive videos that match the query from a massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ without the interference of any video retrieval model.
+For the MVMR task, we further construct practical MVMR benchmarks by extending publicly available VMR datasets.
+Specifically, we categorize positive and negative videos for a specific query from the whole video set of a VMR dataset to construct a massive video set.
+The most critical challenge in converting a VMR dataset into an MVMR dataset is accurately categorizing positive and negative videos for each query.
+However, addressing this issue by manually annotating a large number of videos for each query to determine whether they are positive or negative is exceedingly labor-intensive.
+Therefore, we propose a new automated framework to construct MVMR datasets by employing textual and visual semantic distance verification methods to categorize positive and negative videos for each query, as shown in Figure 2.
+Specifically, we examine the distance between a target query and all videos in a VMR dataset using state-of-the-art text embedding model (Gao et al., 2021) and video-captioning evaluation model (Shi et al., 2022) to define the positive and negative massive video sets.
+Consequently, we construct three MVMR datasets, and we confirm that the introduced datasets contain only 1.5%, 0.2%, and 3.5% falsely defined videos from human evaluation, showing that our approach successfully constructs practical benchmarks.
+On these benchmarks, we evaluate strong baseline VMR models, including state-of-the-art models, and demonstrate that significant overconfidence issues arise in the new scenario (MVMR).
+For example, the state-of-the-art model, MMN, shows an average 35.9% degradation in the performance (R1@0.5 and R5@0.5) for all MVMR datasets, revealing the difficulty of the MVMR task.
+In addition, to enhance the faithfulness of VMR models, we introduce a novel informative sample-weighted learning, CroCs, which stands for Cross-directional informative sample-weighted Contrastive Learning.
+Specifically, CroCs adopts two training mechanisms: (1) weakly-supervised potential negative learning and (2) cross-directional hard-negative learning.
+The weakly-supervised potential negative learning assesses the distance between queries to exclude false-negatives.
+The cross-directional hard-negative learning measures the relevance score between queries and moments predicted by a model to exclude easy negative samples.
+We demonstrate that our CroCs outperforms strong baselines significantly in the MVMR setting, and this result proves that informative sample-weighted constrastive learning successfully mitigates the misinformation overconfidence of a VMR model.
+Moreover, we show that CroCs is also effective in the realistic scenario by adopting the widely-used pipeline, which includes a video retrieval model.
+In summary, this work makes the following contributions:
+  • We propose a new task called MVMR, and construct three MVMR datasets to evaluate the faithfulness of VMR models.
+  • We propose a novel automated dataset construction framework, the first to consider how to construct a trustful massive video retrieval pool via semantic distance verification.
+  • We demonstrate that existing VMR models are significantly vulnerable to misinformation; thus, we introduce a novel informative sample-weighted learning method, CroCs, to enhance the faithfulness of VMR models against misinformation.
+Figure 3.
+Examples of constructed MVMR datasets.
+We visualize positive and negative video sets for a query of the constructed three MVMR datasets.
+A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive.
+The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
+2. Video Moment Retrieval
+Video Moment Retrieval (VMR), which aims to detect a specific moment in a video that matches a given text query, is one of the most challenging problems in the video-language multi-modal domain.
+Therefore, the VMR task has been explored extensively using numerous effective methods  (Zhang et al., 2020a, b; Nan et al., 2021; Gao and Xu, 2021; Liu et al., 2021; Wang et al., 2022; Li et al., 2023).
+Some novel approaches have regarded the problem as question-answering (QA) tasks (Seo et al., 2018; Joshi et al., 2020), by leveraging a QA model to encode a multi-modal representation and then predicting the frames that correspond to the start and end of the moment that match a query (Ghosh et al., 2019; Zhang et al., 2020b).
+However, VMR studies operate within an impractical framework, focusing on searching moments within a given video.
+Consequently, their settings are limited in retrieving relevant information across all media repositories and have not evaluated the faithfulness of VMR models on misinformation.
+Existing studies (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022) have extended the VMR to search on a multiple video pool, naming the task as the Video Corpus Moment Retrieval (VCMR), which focuses on retrieving a moment from multiple videos for a given query.
+The VCMR task considers the most practical usage scenario of VMR.
+However, it is still limited by its inability to evaluate the faithfulness of VMR models in a misinformative context.
+Existing VCMR studies have utilized a framework evaluating moment retrieval performance by integrating video retrieval with VMR; thus, the evaluated performance in the VCMR setting may overwhelmingly depend on video retrieval capability.
+In addition, the VCMR framework simply extends the VMR task by categorizing only a paired positive video moment for a target query as positive while labeling all unpaired videos as negative videos (Yoon et al., 2023; Zhang et al., 2021).
+This approach deviates from the practical scenario of the video moment search since multiple positive videos can be present in whole media collections for a target query.
+Moreover, existing VCMR studies have randomly selected negative samples from all unpaired videos; thus, they have overlooked the possibility of false-negatives when constructing the negative video set for a query, which raises significant trustfulness concerns.
+Some studies (Yoon et al., 2022; Hao et al., 2022; Yoon et al., 2023) have attempted to address unusual behaviors of VMR models, such as bias or spurious correlation, which arise during the training for the VMR and VCMR tasks.
+Yoon et al.
+(2022) has categorized good and bad biases to train a model using only a good bias.
+Hao et al.
+(2022) has mitigated a temporal bias by training a model with temporally shuffled moments.
+However, existing studies have never addressed the overconfidence issue in a misinformative context; thus, it remains a challenging task to solve.
+3. MVMR Problem Definition
+The MVMR task evaluates VMR models under a massive video pool with $k$ positive videos and $n-k$ negative videos for each query $q$.
+Given a positive video set $V^{+}_{q}=\{v^{+}_{1},...,v^{+}_{k}\}$ and a negative video set $V^{-}_{q}=\{v^{-}_{1},...,v^{-}_{n-k}\}$, the MVMR task aims to search for a moment $m=(x_{s},x_{e})$ of a specific video $v$ that matches the query from massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$, where $x_{s}$ and $x_{e}$ mean start and end points of the moment, respectively.
+In contrast to the VCMR task, we only utilize a VMR model $\mathcal{P}$ by inputting a target query $q$ and each video $v$ in $V^{+,-}_{q}$ to compute the matching score $p=\mathcal{P}(m|q,v)$ for each moment $m$.
+We finally determine the optimal moment $m^{*}=(x^{*}_{s},x^{*}_{e})$ by comparing the matching score of all moments in $V^{+,-}_{q}$ as $m^{*}=argmax_{m}\mathcal{P}(m|q,v)$.
           Dataset      # Queries / # Videos             Avg Len (sec)             Avg # Moments             Avg            # Max Positive /
@@
 $\text{MVMR}_{\text{TACoS}}$ 2055 / 25 28.55 / 367.15 2.24 7.31 5 / 5
-Table 1: Summary of datasets.
-Remarkably, the queries of $\text{MVMR}_{\text{Charades-STA}}$ and $\text{MVMR}_{\text{TACoS}}$ generally include multiple positive moments, and it proves that existing VCMR studies are exposed to the risk of false negatives.
-2.2 Video Corpus Moment Retrieval
-Existing studies (Escorcia et al., 2019; Lei et al., 2020; Ma and Ngo, 2022; Jung et al., 2022) have tried to extend the NLVL to search on multiple video pools, naming it the Video Corpus Moment Retrieval (VCMR) task, which focuses on retrieving a moment from multiple videos for a given query.
-However, the VCMR simply extends the NLVL task by designating only one matched positive video moment for a target query as a positive, while labeling all unmatched videos as negative videos.
-This approach deviates from the practical scenario of the video moment search since multiple positive videos can be present for a target query.
-Moreover, existing VCMR studies have failed to consider the possibility of false-negative when constructing the negative video set for a specific query, which poses a significant reliability issue.
-3 Problem Definition
-MVMR task evaluates video moment retrieval models under an MVMR dataset with $k$ positive videos and $n-k$ negative videos for each query.
-Given a positive video set $V^{+}_{q}=\{v^{+}_{1},...,v^{+}_{k}\}$ and a negative video set $V^{-}_{q}=\{v^{-}_{1},...,v^{-}_{n-k}\}$, the MVMR task aims to localize a moment $(x_{s},x_{e})$ of a specific video $v$ that matches the query from massive video set $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$, where $x_{s}$ and $x_{e}$ mean start and end points of the moment, respectively.
-In MVMR task, we aim to retrieve an optimal positive moment $(x^{*}_{s},x^{*}_{e})$ that matches semantically with the text query $q$.
-We derive confidence scores of moments for the whole $n$ massive videos and select the moment with the highest score as a prediction.
-4 MVMR Dataset Construction
-Our MVMR setting automatically reconstructs an NLVL dataset by expanding it to associate each query with multiple positive and negative moments.
-The paired positive and negative videos are carefully selected with semantic similarity check techniques.
-In the following, we detail our process for constructing an MVMR dataset and present the analyses of the three constructed MVMR datasets.
-4.1 Semantic Similarity Check
-This section examines the similarity between a target query $q$ and a video $v$ using semantic text embedding and video captioning evaluation models.
-Query Similarity-based Check.
-NLVL datasets consist of video set $V=\{v_{1},...,v_{l}\}$ and paired query set $Q_{v}=\{q_{1},...,q_{m}\}$ for each video $v$, where $l$ and $m$ is the number of videos and text queries included in a dataset, respectively.
+Table 1.
+Summary of datasets.
+Remarkably, the queries of $\text{MVMR}_{\text{Charades-STA}}$ and $\text{MVMR}_{\text{TACoS}}$ generally include multiple positive moments, and it empirically proves that a massive video set construction using random sampling is inadequate to construct the MVMR dataset since it is exposed to the risk of false negatives.
+4. MVMR Dataset Construction
+Our MVMR setting automatically reconstructs a VMR dataset by expanding it to associate each query with multiple positive and negative moments.
+Each VMR dataset consists of multiple videos with some query-moment pairs, and our method carefully categorizes the positive and negative sets using two semantic distance verification techniques: Query Distance-based Verification and Query-video Distance-based Verification.
+In the following, we detail our process for constructing the MVMR benchmark and present the analyses of three constructed MVMR datasets.
+4.1. Semantic Distance Verification
+This section describes the verification process of the distance between a target query $q$ and a video $v$ using semantic text embedding and video captioning evaluation models.
+Query Distance-based Verification.
+VMR datasets consist of video set $V=\{v_{1},...,v_{l}\}$ and paired query set $Q_{v}=\{q_{1},...,q_{m}\}$ for each video $v$, where $l$ and $m$ is the number of videos and text queries included in a dataset, respectively.
 Since each text query semantically describes the paired moment, we use queries to obtain semantic information about the paired video.
-For example, suppose that a video $v$ is paired with a query $q$, "The person gets out a knife", in a moment $(x_{s},x_{e})$.
-Since a target query $\hat{q}$, "The person takes out a knife", is highly similar to the query $q$, $\hat{q}$ is associated with the moment $(x_{s},x_{e})$ of the video $v$.
-We use a semantic text embedding model, SimCSE Gao et al.
-(2021), to distinguish positive and negative videos for each query.
-Specifically, we first obtain embeddings $e_{\hat{q}}$ and $e_{q_{1}},...,e_{q_{m}}$ for the target text query $\hat{q}$ and all video text queries $q_{1},...,q_{m}$ using SimCSE, respectively.
-We calculate cosine similarity $s_{te}(\hat{q},q_{i})$ between $e_{\hat{q}}$ and $e_{q_{i}}$ to quantify the semantic matching, and select the maximum similarity among $s_{te}(\hat{q},q_{1}),...,s_{te}(\hat{q},q_{m})$ to use it as the query-video matching score $s_{te}(\hat{q},v)$ of the target query $\hat{q}$ and the video $v$.
+For example, suppose that a video $v$ is paired with a query $q$, ”The person gets out a knife”, in a moment $(x_{s},x_{e})$.
+Since a target query $\hat{q}$, ”The person takes out a knife”, is highly similar to the query $q$, $\hat{q}$ is associated with the moment $(x_{s},x_{e})$ of the video $v$.
+We use a semantic text embedding model, SimCSE (Gao et al., 2021), to distinguish positive and negative videos for each query.
+Specifically, we first obtain embeddings $e_{\hat{q}}$ and $\{e_{q_{1}},...,e_{q_{m}}\}$ for the target text query $\hat{q}$ and all video text queries $\{q_{1},...,q_{m}\}$ using SimCSE, respectively.
+Then, we calculate cosine similarity $s_{te}(\hat{q},q_{i})$ between $e_{\hat{q}}$ and $e_{q_{i}}$ to quantify the semantic matching, and select the maximum similarity among $\{s_{te}(\hat{q},q_{1}),...,s_{te}(\hat{q},q_{m})\}$ to use it as the query-video matching score $s_{te}(\hat{q},v)$ of the target query $\hat{q}$ and the video $v$.
 We select max aggregation to obtain $s_{te}(\hat{q},v)$ since the video should be regarded as a positive video, even if only one similar query to a target query exists in a video.
-Figure 2: Examples of constructed MVMR datasets.
-We visualize positive and negative video sets for the query “person closes the door" of the $\text{MVMR}_{\text{Charades-STA}}$.
-A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive.
-The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
-Query-video Similarity-based Check.
-Queries associated with a video may describe the video limitedly since they may not cover all of its context.
-Therefore, we examine a visual semantic matching between a query and a video to filter positive and negative videos.
-We use EMScore Shi et al.
-(2022), a video captioning evaluation model, to quantify the visual semantic similarity $s_{tv}$, used as a complementary filtering method on videos primarily filtered through the query similarity $s_{te}$.
-4.2 Distinguishing Positive and Negative
-This section describes constructing positive and negative video sets using the calculated query-video similarity, $s_{te}$ and $s_{tv}$.
-Positive candidates are distinguished by primarily excluding videos $s_{te}(q,v)<t_{te}^{+}$, where $t_{te}^{+}$ is a hyper-parameter.
+Query-video Distance-based Verification.
+Paired queries may describe the video limitedly since they may not cover all its context.
+We examine a visual semantic matching between a query and a video to categorize positive and negative videos.
+We use EMScore (Shi et al., 2022), a video captioning evaluation model, to quantify the visual similarity $s_{tv}(q,v)$, used as a complementary filtering method on videos primarily selected through the query similarity $s_{te}(q,v)$.
+4.2. Categorizing Positive and Negative sets
+This section describes constructing positive and negative video sets using the calculated query-video similarities, $s_{te}$ and $s_{tv}$.
+Positive candidates are identified by primarily excluding videos with $s_{te}(q,v)<t_{te}^{+}$, where $t_{te}^{+}$ is a hyper-parameter.
 We further exclude videos that have lower similarity than the mean $s_{tv}$ of all golden positive query-video pairs $(q,v^{*})$ for all queries.
-We construct final MVMR positives by randomly sampling $k$ videos from this filtered set of positive candidates.
-Negative candidates are defined by excluding videos with $s_{te}(q,v)>t_{te}^{-}$, where $t_{te}^{-}$ is also a hyper-parameter.
-We additionally exclude videos that have higher similarity than the mean $s_{tv}$ of query-video pairs $(q,v^{-})$, which are primarily filtered as negatives using $s_{te}$.
+We construct final MVMR positives by randomly sampling $k$ videos from this categorized set of positive candidates.
+Negative candidates are identified by excluding videos with $s_{te}(q,v)>t_{te}^{-}$, where $t_{te}^{-}$ is also a hyper-parameter.
+We additionally exclude videos that have higher similarity than the mean $s_{tv}$ of query-video pairs $(q,v^{-})$, which are primarily categorized as negatives using $s_{te}$.
 Like MVMR positives, we construct final MVMR negatives by randomly sampling $(n-k)$ videos from this set of negative candidates.
-From this process, a total of $n$ videos retrieval pool $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ is obtained for each query.
-5 MVMR Dataset Analysis
-5.1 MVMR Settings.
-This section describes the used NLVL datasets and detailed settings to derive MVMR positive and negative candidates using SimCSE and EMScore.
-NLVL Datasets.
-We extend widely-used NLVL datasets (Charades-STA Gao et al.
-(2017), ActivityNet Krishna et al.
-(2017), and TACoS Regneri et al.
-(2013)) to construct three MVMR evaluation datasets, and name the constructed datasets as $\text{MVMR}_{\textsc{\{\$datasource\}}}$.
-Each NLVL dataset consists of multiple videos with query-moment pairs.
-The details of each NLVL dataset are shown in Appendix A.
+From this process, a total of $n$ video retrieval pool $V^{+,-}_{q}=\{v^{+}_{1},...,v^{+}_{k},v^{-}_{1},...,v^{-}_{n-k}\}$ is obtained for each query.
+5. MVMR Dataset Analysis
+5.1. MVMR Settings.
+This section describes the utilized VMR datasets for our MVMR framework and detailed settings to derive MVMR positive and negative candidates using SimCSE and EMScore.
+VMR Datasets.
+We extend widely-used VMR datasets, Charades-STA (Gao et al., 2017), ActivityNet (Krishna et al., 2017), and TACoS (Regneri et al., 2013), to construct three MVMR benchmarks.
+We name the constructed MVMR datasets as $\text{MVMR}_{\textsc{\{\$datasource\}}}$.
 Filtering Settings.
-We use STS16 dataset Agirre et al.
-(2016) to empirically validate varying SimCSE filtering hyper-parameters, $t^{+}_{te}$ and $t^{-}_{te}$.
-The dataset provides text pairs and human-annotated similarity scores (five-point Likert scale); two texts are regarded as similar when the score is higher than 3.
+We use the STS16 dataset (Agirre et al., 2016) to empirically validate varying SimCSE filtering hyper-parameters, $t^{+}_{te}$ and $t^{-}_{te}$.
+The dataset provides text pairs and human-annotated similarity scores (five-point Likert scale), and two texts are considered similar when the score exceeds 3.
 We analyze the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ to derive MVMR positive candidates since it is regarded as the best threshold to distinguish similar queries.
-We also select $t^{-}_{te}=0.5$ to filter MVMR negative candidates.
-The distribution of SimCSE and human-annotated scores are shown in Appendix B.
-Figure 3: Qualitative Analysis for Filtering Methods.
+Similarly, we select $t^{-}_{te}=0.5$ to identify MVMR negative candidates.
+The distribution of SimCSE and human-annotated scores are shown in Appendix A.
+Figure 4.
+Qualitative Analysis for Filtering Methods.
 We visualize the derived similarity scores of SimCSE and EMScore to verify the constructed MVMR datasets.
 We use T-SNE to reduce the dimension of each query embedding for displaying each query (dot) of SimCSE Similarity Distribution.
-5.2 MVMR Analysis
-We set a video retrieval pool as $n$, containing at most $k$ positive videos for each query.
+5.2. MVMR Analysis
+We fixedly set the size of the massive video pool as $n$, containing at most $k$ positive videos for each query.
 If the number of MVMR positive candidates derived for each query is less than $k$, we include all positive candidates as positive samples.
@@
 Datasets Statistic.
-We set the number of videos in the $\text{MVMR}_{\text{Charades-STA}}$ and the $\text{MVMR}_{\text{ActivityNet}}$ retrieval pools as $n=50$, and the maximum number of positive moments per query $k$ is set to $5$.
+We determine the number of videos in the $\text{MVMR}_{\text{Charades-STA}}$ and the $\text{MVMR}_{\text{ActivityNet}}$ retrieval pools as $n=50$, and the maximum number of positive moments per query $k$ is set to $5$.
 The former consists of 3,716 queries for 1,334 videos, each containing an average of 3.07 positive moments.
 The latter contains 16,941 queries for 4,885 videos, each including an average of 1.11 positive videos.
-We set the number of videos for the $\text{MVMR}_{\text{TACoS}}$ and the maximum number of positive moments per query as $5$ and $5$, respectively, since TACoS test set contains only 25 videos.
+For the $\text{MVMR}_{\text{TACoS}}$ dataset, We set the size of the video pool and the maximum number of positive moments per query as $5$ and $5$, respectively, since the TACoS test set contains only 25 videos totally.
 It includes 2,055 queries for 25 videos, each containing an average of 2.24 positive videos.
-The summary of the MVMR datasets is shown in Table 1.
+The summary of the utilized VMR datasets and the constructed MVMR datasets is shown in Table 1.
 Dataset Quality.
 […]
-[…]
-Figure 4: Cohesiveness of the Positive and Negative sets.
-The X and Y axes mean the word co-occurrence ratio and its frequency, respectively.
-The closer the word’s co-occurrence ratio approaches 1.0, the more cohesive it is.
-Human Evaluation.
-We recruit crowd workers and ask them to examine 100 queries manually and their paired videos (by our method) and answer whether they are correctly labeled.
-Overall, we analyze 4,900, 4,900, and 400 videos for Charades-STA, ActivityNet, and TACoS, respectively, since each MVMR dataset consists of 49, 49, and 4 newly added videos (either positive or negative samples).
-The original video (i.e., golden positive sample) is excluded from this investigation.
-From human evaluation, we confirm that the newly introduced dataset contains only 1.5%, 0.2%, and 3.5% falsely categorized videos for the three datasets, respectively.
-TACoS exhibits a relatively higher query-relevant video ratio than the other two datasets since its test set includes only 25 videos and consists of considerably similar cooking activities within the kitchen.
-Analysis on Semantic Filtering Methods.
-We visualize the scores derived using SimCSE and EMScore to analyze our MVMR datasets qualitatively.
-Figure 3-(a) displays SimCSE embeddings of all queries in each dataset, illustrating the similarity between all queries and a specific target query, "person turn a light on." (left) and "He is using a push mower to mow the grass." (right).
-The figure of TACoS can also be found in Appendix B.
-The figures show that queries exhibit well-defined clusters for Charades-STA and TACoS, revealing the effectiveness of our query similarity-based filtering.
+Analysis on Semantic Distance Verification.
+We visualize the results derived using SimCSE and EMScore to analyze our MVMR datasets qualitatively.
+Figure 4-(a) displays SimCSE embeddings of all queries in each dataset, illustrating the similarity between all queries and a specific target query, ”person turn a light on.” (left) and ”He is using a push mower to mow the grass.” (right).
+The figure of TACoS can also be found in Appendix A.
+The figures show that queries exhibit well-defined clusters for Charades-STA and TACoS, revealing the effectiveness of our Query Distance-based filtering.
 These results are also supported by Figure 4-(b) SimCSE score histogram, showing that Charades-STA and TACoS have many similar queries.
@@
 The red and blue histograms correspond to the EMScore distribution of positive and negative samples, respectively.
-For the ActivityNet, EMScore effectively distinguishes positives and negatives.
+In the case of ActivityNet, EMScore effectively distinguishes positives and negatives.
 This success is attributed to that ActivityNet includes detailed queries and videos with diverse features.
-According to the findings presented in this section, the filtering method’s efficacy varies depending on the dataset’s characteristics.
-Consequently, a combination of two filtering methods (i.g., SimCSE and EMScore) should be employed in a complementary manner to ensure the construction of a reliable MVMR dataset.
-6 Methods: CroCs
-Figure 5: CroCs Overview.
-We adopt informative negative contrastive learning to solve the MVMR task.
+According to the findings presented in this section, the Semantic Distance Verification method’s efficacy varies depending on the dataset’s characteristics.
+Consequently, a combination of two filtering methods (i.g., SimCSE and EMScore) should be employed in a complementary manner to ensure the construction of a trustful MVMR dataset.
+Human Evaluation.
+We recruit crowd workers fluent in English through the university’s online community.
+We ask them to examine 100 queries manually and their paired videos (by our method) and answer whether they are correctly labeled.
+Overall, we analyze 4,900, 4,900, and 400 videos for Charades-STA, ActivityNet, and TACoS, respectively, since each MVMR dataset consists of 49, 49, and 4 newly added videos (either positive or negative samples).
+The original videos (i.e., golden positive samples) are excluded from this investigation.
+Based on human evaluation, we confirm that the newly introduced dataset contains only 1.5%, 0.2%, and 3.5% falsely categorized videos for the three datasets, respectively.
+TACoS exhibits a relatively higher query-relevant video ratio than the other datasets since its test set includes only 25 videos and consists of significantly similar cooking activities within the kitchen.
+6. Methods: CroCs
+Figure 5.
+CroCs Overview.
+We adopt the informative sample-weighted mutual matching learning to solve the MVMR task.
 The dots and triangles are the features of moments and texts.
-The blue dash lines are matched moment-text pairs to be pulled in, while the red dash lines are negative samples of intra/inter-video to be pushed away.
+The blue dash line is matched moment-text pairs to be pulled in, while the red dash lines are negative samples of intra/inter-video to be pushed away.
 The yellow and orange dash lines are unmatched moment-text pairs, but not to train by filtering out since they are easy and false-negatives.
-This section introduces a novel contrastive learning method, CroCs, which stands for Cross-directional Hard and Reliable Negative Contrastive Learning, to solve the MVMR task.
-A key challenge in solving the MVMR task is distinguishing positives from multiple negative distractors.
-Therefore, we select MMN Wang et al.
-(2022) since it is the optimal baseline model for the MVMR task as it uses contrastive learning to distinguish negative samples from positives effectively.
-We use the MMN as the backbone of our model and enhance its ability by applying CroCs in training procedures to solve the challenge in the MVMR task.
+To solve the MVMR task, this section introduces a novel informative sample-weighted learning method, CroCs, which stands for Cross-directional informative sample-weighted Contrastive Learning.
+A key challenge in solving the MVMR task is distinguishing the representation of positives from multiple negative distractors.
+Therefore, we select MMN (Wang et al., 2022) as the backbone of our model since it is the optimal baseline for the MVMR task since it uses contrastive learning to distinguish negative samples from positives effectively.
+On the MMN architecture, we apply CroCs in contrastive learning procedures to solve the challenge in the MVMR task.
 MMN has limitations in neglecting the potential presence of false-negative video moments since it regards all randomly selected in-batch samples as negatives in contrastive learning.
 Furthermore, it overlooks the importance gap among negative samples.
-Our CroCs solves these two limitations of the MMN by adopting two contrastive learning procedures: (1) Weakly-supervised potential negative learning and (2) Informative hard-negative learning.
-6.1 Weakly-supervised Contrastive Learning
-A critical consideration in the contrastive learning framework is the potential presence of false-negatives.
-This issue arises when all videos, except for the golden video, are categorized as negative training samples.
+Our CroCs solves these two limitations of MMN by adopting two training mechanisms: (1) Weakly-supervised potential negative learning and (2) Cross-directional hard-negative learning.
+6.1. Weakly-supervised Potential Negative Learning
+MMN randomly selects negative samples from the whole samples for contrastive learning.
+This random sampling process has a significant limitation due to the potential presence of false-negatives since it categorizes all samples, except for the golden one, as negative training samples.
 To mitigate this concern, we propose excluding false-negative videos to refine the contrastive learning process.
-Specifically, we adopt the SimCSE filtering to identify and exclude false-negative videos as follows:
-$V^{-}_{tr}(q)=\{v|s_{te}(q,v)<t_{tr}\},\>\>\ \forall{v}\in V\>$ (1)
+Specifically, we adopt a potential negative filtering method to identify and exclude false-negative videos as follows:
+(1) $V^{-}_{tr}(q)=\{v|s_{te}(q,v)<t_{tr}\},\>\>\ \forall{v}\in V$
 where $V^{-}_{tr}(q)$ is a negative sample set of a specific query $q$ to use during the contrastive learning procedure.
-$s_{te}(q,v)$ is query-video similarity described in Section 4.1.
-$t_{tr}$ is a hyper-parameter for SimCSE score filtering.
-If a specific video shows a value of $s_{te}(q,v)$ above the threshold $t_{tr}$ for target query $q$, we classify $v$ as false-negative for $q$.
+$s_{te}(q,v)$ is the query-video similarity described in Section 4.1.
+$t_{tr}$ is a hyper-parameter for the SimCSE score filtering.
+If a specific video $v$ shows a value of $s_{te}(q,v)$ above the threshold $t_{tr}$ for a target query $q$, we classify $v$ as false-negative for $q$.
 Therefore, we exclude $v$ in negative samples of contrastive learning for $q$.
-6.2 Hard-negatives Contrastive Learning
-Easy or false-negatives may distract a model from learning proper knowledge during the training process.
-Inspired by recent progress in using hard-negative sampling for the document retrieval task Zhou et al.
-(2022), we categorize negatives according to the matching score between a target query and negative moments as follows: (1) Negatives that are clearly irrelevant and have low matching scores should be sampled less frequently (Easy-negative); (2) Negatives that are highly relevant and have high matching scores should also be sampled less frequently (False-negative); (3) Negatives that are uncertain and have matching scores similar to true-positives should be sampled more frequently since they provide useful information (Hard-negative).
-Based on these criteria, we distinguish hard-negative samples and adopt a retraining process for a model already trained on a specific dataset using only hard-negative samples.
-We define the negative moments and query sampling distribution as follows:
-$p^{\prime}(m|q)\propto\exp{(-(r(q,m)-\overline{r}(q,m^{+}))^{2})},\>\ \forall m\in M^{-}_{q}$ (2)
-$p^{\prime}(q|m)\propto\exp{(-(r(q,m)-\overline{r}(q^{+},m))^{2})},\>\ \forall q\in Q^{-}_{m}\>\>$
-where $r(q,m)$ is a matching score between a query and a moment calculated using the already trained video moment retrieval model, and $\overline{r}(q,m^{+})$ and $\overline{r}(q^{+},m)$ is mean matching scores for all golden positive moments and queries, respectively.
+6.2. Cross-directional Hard-negative Learning
+Easy or false-negative samples may distract a model from learning proper knowledge during the training procedure.
+Inspired by recent progress in using hard-negative sampling for the document retrieval task (Zhou et al., 2022), we classify negative samples according to the matching score between a target query and negative moments as follows: (1) Negative samples that are clearly irrelevant and have low matching scores should be sampled less frequently (Easy-negative); (2) Negative samples that are highly relevant and have high matching scores should also be sampled less frequently (False-negative); (3) Negative samples that are uncertain and have matching scores similar to true-positives should be sampled more frequently since they provide useful information (Hard-negative).
+Based on these criteria, we distinguish informative hard-negative samples and adopt a retraining process for a model pre-trained on a specific dataset, using only hard-negative samples.
+We define the negative sampling distribution for the moment and query as follows:
+(2) $p^{\prime}(m|q)\propto\exp{(-(r(q,m)-\overline{r}(q,m^{+}))^{2})},\>\ \forall m\in M^{-}_{q}$
+$p^{\prime}(q|m)\propto\exp{(-(r(q,m)-\overline{r}(q^{+},m))^{2})},\>\ \forall q\in Q^{-}_{m}$
+where $r(q,m)$ is a matching score between a query and a moment calculated using a pre-trained video moment retrieval model, and $\overline{r}(q,m^{+})$ and $\overline{r}(q^{+},m)$ is mean matching scores for all golden positive moments and queries, respectively.
 $M^{-}_{q}$ and $Q^{-}_{m}$ is potential negative candidates for a query $q$ and a moment $m$, respectively.
-$p\textquoteright(m|q)$ means the probability that moment $m$ is sampled as negative when given query $q$.
-Therefore, negative candidate moments with a close distance to the average matching score of golden query-moment pairs have a high sampling probability.
-Likewise, $p\textquoteright(q|m)$ means the probability that query $q$ is sampled as negative when given moment $m$.
-We also reformulate $p\textquoteright(m|q)$ to $p\textquoteright(v|q)$ for computational efficiency in a sampling procedure as follows:
-$r(q,v)=\max{(\{r(q,m_{1}),...,r(q,m_{l})\})},\>\>\ \forall m_{i}\in v\>\>\>\>\>\>$ (3)
-$p^{\prime}(v|q)\propto\exp{(-(r(q,v)-\overline{r}(q,v^{+}))^{2})},\>\>\ \forall v\in V^{-}_{q}\>\>\>\>\>\>$
-$p^{\prime}(q|v)$ can be formulated similarly by calculating $\overline{r}$ about positive queries for each video.
-6.3 Reliable Mutual Matching Network
-This section describes a model architecture to apply the proposed contrastive learning method, CroCs.
+$p’(m|q)$ means the probability that a moment $m$ is sampled as negative when given a query $q$.
+Specifically, $p’(m|q)$ is computed by the distance between $r(q,m)$ and $\overline{r}(q,m^{+})$; thus, the closer the distance, the more frequently sampled as negative since a closer sample is a hard-negative sample not easy to distinguish by a model.
+In other words, negative candidate moments near the average matching score of golden query-moment pairs have a high sampling probability.
+Likewise, $p’(q|m)$ means the probability that a query $q$ is sampled as negative when given a moment $m$.
+We also reformulate $p’(m|q)$ to $p’(v|q)$ for computational efficiency in a sampling procedure as follows:
+(3) $r(q,v)=\max{(\{r(q,m_{1}),...,r(q,m_{l})\})},\>\>\ \forall m_{i}\in v$
+$p^{\prime}(v|q)\propto\exp{(-(r(q,v)-\overline{r}(q,v^{+}))^{2})},\>\>\ \forall v\in V^{-}_{q}$
+$p^{\prime}(q|v)$ can be formulated similarly by calculating $\overline{r}$ for positive queries about each video.
+6.3. Informative Sample-weighted Mutual Matching Learning
+This section describes an informative sample-weighted contrastive learning method, CroCs.
+Model architecture.
 We adopt a bi-encoder architecture with a late modality fusion by an inner product in the joint visual-text representational space following the baseline MMN.
-By adopting a bi-encoder architecture, we can efficiently search for moments with pre-calculated video moment representations given a user query.
+By adopting a bi-encoder architecture, we can efficiently search for moments with pre-computed video moment representations when given a user query.
 Our whole model architecture is shown in Figure 5.
 Our model represents each query and video moment using query and video encoders to two joint visual-text spaces (IoU and mutual matching spaces).
-From the joint visual-text space, we compute the matching score between the query and video moment representations and select the best video moment.
-The details of the model architecture are described in the appendix C.1.
-We select the binary cross entropy and contrastive learning losses and train our model following two steps: (1) the first training step using the two loss functions (binary cross entropy and contrastive learning losses) filtering out false-negatives described in Section 6.1; (2) the second training step using two loss functions with only hard-negatives described in Section 6.2.
-To calculate the similarity between the text and video moment features projected in the IoU space, we compute the cosine similarity $s^{iou}$.
-To adjust the range of the final prediction, we multiply the cosine similarity by a factor of 10, following the MMN.
+From the joint visual-text spaces, we compute the matching score between query and video moment representations and select the best video moment.
+The details of the model architecture are described in the appendix B.1.
+Informative Sample-weighted Mutual Matching.
+We select the binary cross entropy and contrastive learning losses and train our model following two steps: (1) the first training step using the two loss functions (binary cross entropy and contrastive learning losses) filtering out potential false-negatives described in Section 6.1; (2) the second training step using two loss functions with only informative hard-negatives described in Section 6.2.
+To calculate the similarity between the text and video moment representations projected in the IoU space, we compute the cosine similarity $s^{iou}$.
+To adjust the range of the final prediction, we multiply the cosine similarity by a factor of 10, following MMN.
 This amplification results in the final prediction $p^{iou}_{i}=\sigma(10\cdot s^{iou}_{i})$, where $\sigma$ represents the sigmoid function.
 The binary cross-entropy loss is then calculated as follows:
-$L_{bce}=-\frac{1}{C}\sum^{C}_{i=1}{(y_{i}\log{p^{iou}_{i}}+(1-y_{i})\log{(1-p^{iou}_{i})})}$ (4)
-where $p^{iou}_{i}$ is the confidence score of each moment, and $C$ is the total number of valid candidates. We calculate $L_{bce}$ for positive and true-negative samples derived in Section 6.1.
+(4) $\mathcal{L}_{bce}=-\frac{1}{C}\sum^{C}_{i=1}{(y_{i}\log{p^{iou}_{i}}+(1-y_{i})\log{(1-p^{iou}_{i})})}$
+where $p^{iou}_{i}$ is the confidence score of each moment, and $C$ is the total number of valid candidates.
+We calculate $\mathcal{L}_{bce}$ for positive and potential true-negative samples derived in Section 6.1.
 We use the cross-directional contrastive learning loss to effectively train the model using the informative negatives derived from the methods described in Section 6.1 and 6.2.
-The mutual matching contrastive learning loss considering the informative negative samples selection process is as follows:
-$p^{*}(q_{i}|m)=\frac{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})}{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})+\sum_{j\in\mathcal{I}_{m}^{-}}{\exp(f_{j}^{q\mathsf{T}}f^{m}/\tau_{m})}}\>\>\>$ (5)
-$p^{*}(m_{i}|q)=\frac{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})}{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})+\sum_{j\in\mathcal{I}_{q}^{-}}{\exp(f_{j}^{m\mathsf{T}}f^{q}/\tau_{q})}}\>\>\>\>\>$
-$L_{rmm}=-(\sum^{N}_{i=1}{(\log{p^{*}(i_{m}|q_{i})}})+\sum^{N}_{i=1}{(\log{p^{*}(i_{q}|m_{i})}}))\>\>\>\>\>\>\>\>\>\>\>$
-Dataset Model R@1 R@1 R@1 R@5 R@5 R@5 R@20 R@20 R@20 R@50 R@50 R@50
-IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7
-$\text{MVMR}_{\text{Charades-STA}}$ CET 3.96 2.96 2.02 15.34 11.79 6.97 39.85 32.21 19.91 59.58 50.54 33.88
-BET 4.57 2.99 1.94 14.26 10.98 7.00 31.05 25.11 16.55 45.88 37.86 26.08
-G2L 4.30 2.78 1.97 14.34 10.59 7.36 40.89 34.74 21.33 71.92 62.72 44.82
-MMN 13.99 12.78 8.40 37.86 33.29 22.36 64.80 57.83 41.71 80.11 73.09 56.75
-${}^{\dagger}$ 18.38 16.31 10.76 43.86 39.05 26.99 69.48 62.89 46.39 82.88 76.08 60.28
-CroCs 19.27 17.38 11.44 43.54 39.02 27.91 68.08 60.98 46.39 82.99 76.29 60.76
-$\text{MVMR}_{\text{ActivityNet}}$ CET 1.26 0.90 0.56 2.81 2.32 1.75 5.41 4.73 3.87 8.97 8.02 6.56
-BET 1.56 1.16 0.76 3.02 2.55 2.03 5.30 4.64 3.80 8.13 7.10 5.88
-G2L 1.66 1.32 0.96 7.82 6.43 4.55 26.84 21.66 14.81 62.99 50.07 32.99
-MMN 13.82 10.71 6.93 31.54 25.04 16.55 54.79 44.50 30.62 73.08 62.09 45.76
-${}^{\dagger}$ 16.46 12.68 7.87 36.03 28.47 18.16 57.91 46.52 31.12 73.82 62.12 44.89
-CroCs 20.63 15.58 9.51 44.13 34.70 22.40 67.01 55.58 38.55 79.91 70.02 53.20
-$\text{MVMR}_{\text{TACoS}}$ CET 8.95 5.79 3.89 26.18 18.88 12.17 54.84 42.53 23.70 - - -
-BET 8.22 5.30 2.97 29.46 14.84 8.32 55.67 39.76 21.51 - - -
-G2L 8.72 5.82 4.31 30.61 21.51 12.29 56.52 42.35 22.52 - - -
-MMN 10.56 8.61 5.60 36.16 25.84 15.47 58.20 43.89 23.80 - - -
-${}^{\dagger}$ 13.43 10.75 6.76 36.45 27.01 15.28 58.30 44.38 25.16 - - -
-CroCs 13.24 10.75 7.30 39.22 29.68 16.50 60.05 44.77 24.62 - - -
-Table 2: Experimental Results on MVMR datasets.
-Bolded and under-lined results indicate the $1^{st}$ and $2^{nd}$ best performance, respectively.
-${}^{\dagger}$ is the model using only weakly-supervised contrastive learning (§6.1).
-CroCs uses both of two contrastive learning methods (§6.1 and §6.2).
-We train all the baselines as three trials and report the averaged MVMR scores.
+The reliable mutual matching contrastive learning loss considering the informative negative samples selection process is as follows:
+(5) $p^{*}(q_{i}|m)=\frac{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})}{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})+\sum_{j\in\mathcal{I}_{m}^{-}}{\exp(f_{j}^{q\mathsf{T}}f^{m}/\tau_{m})}}$
+$p^{*}(m_{i}|q)=\frac{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})}{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})+\sum_{j\in\mathcal{I}_{q}^{-}}{\exp(f_{j}^{m\mathsf{T}}f^{q}/\tau_{q})}}$
+$L_{rmm}=-(\sum^{N}_{i}{\sum^{M}_{j}\log{p^{*}(m_{ij}|q_{i})}}+\sum^{N}_{i}{\sum^{M}_{j}\log{p^{*}(q_{ij}|m_{i})}})$
 where $q_{i}$ and $m_{i}$ are corresponding positive query and moment for each moment and query, respectively.
-$f^{m}$ and $f^{q}$ are the moment and query features in the joint visual-text space.
+$f^{m}$ and $f^{q}$ are moment and query features in the joint visual-text space.
 $\tau_{q}$ and $\tau_{m}$ are temperatures.
 $\mathcal{I}^{-}$ means the indices of informative negatives derived using our sampling methods.
-The final loss, denoted as $L$, is formulated as a linear combination of the binary cross-entropy loss and the reliable mutual matching loss.
-The matching score $s$, for a specific moment given the text query, is obtained by multiplying the iou score $s^{iou}$ with the reliable mutual matching score $s^{rmm}$.
-The reliable mutual matching score is calculated for the features in the mutual matching space in the same way as the iou score.
-$L=L_{bce}+\lambda L_{rmm},\>\>\>\>\ s=s^{iou}\cdot s^{rmm}$ (6)
-7 Experiments
-7.1 Experimental Settings
-This section describes the experimental settings of the CroCs and other baselines. More details are shown in Appendix C.
+The final loss, denoted as $\mathcal{L}$, is formulated as a linear combination of the binary cross-entropy loss and the reliable mutual matching loss.
+The matching score $s$, for a specific moment given a query, is obtained by multiplying the iou score $s^{iou}$ with the reliable mutual matching score $s^{rmm}$.
+The reliable mutual matching score is calculated for features in the mutual matching space in the same way as the iou score.
+(6) $\mathcal{L}=\mathcal{L}_{bce}+\lambda\mathcal{L}_{rmm},\>\>\>\>\ s=s^{iou}\cdot s^{rmm}$
+7. Experiments
+7.1. Experimental Settings
+This section describes the experimental settings of the CroCs and other baselines. More details are shown in Appendix B.
 Implementation Details.
-We use standard off-the-shelf video feature extractors without any fine-tuning: VGG Simonyan and Zisserman (2015) feature for Charades-STA; C3D Tran et al.
-(2015) feature for ActivityNet and TACoS following Wang et al.
-(2022).
-We set the filtering hyper-parameter as $t_{tr}=0.9$ for the weakly-supervised contrastive learning.
+We use standard off-the-shelf video feature extractors without any fine-tuning: VGG (Simonyan and Zisserman, 2015) feature for Charades-STA; C3D (Tran et al., 2015) feature for ActivityNet and TACoS following (Wang et al., 2022).
+We set the filtering hyper-parameter as $t_{tr}=0.9$ for the weakly-supervised potential negative contrastive learning.
 We choose the number of the CroCs hard negative samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
@@
 Baselines.
-We first select the state-of-the-art models, MMN Wang et al.
-(2022) and G2L Li et al.
-(2023), as strong baselines.
-As the MVMR  task can be formulated as an open-domain question answering (QA) task, we further choose two types of QA models that employ a cross-encoder architecture (i.e., Cross-Encoder Transformer) and a dual-encoder architecture (i.e., Bi-Encoder Transformer), respectively.
-These architectures are widely adopted in existing NLVL tasks Zhang et al.
-(2020b); Gao and Xu (2021) due to the closeness of the two tasks.
+We first select the state-of-the-art models, MMN (Wang et al., 2022) and G2L (Li et al., 2023), as strong baselines.
+As the MVMR  task can be formulated as an open-domain question answering (QA) task, we further choose two types of QA models that employ a cross-encoder architecture, Cross-Encoder Transformer (CET), and a dual-encoder architecture, Bi-Encoder Transformer (BET).
+These architectures are widely adopted in existing VMR tasks due to the closeness of the two tasks.
 Evaluation Metrics.
@@
 It is defined as the percentage of queries having at least one correctly retrieved moment, i.e., IoU $\geq m$, in the top-$n$ derived moments.
-The existing NLVL studies have reported the results of $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5\}$ Wang et al.
-(2022).
-But, since our MVMR setting searches for massive videos, we report the extended results as $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20,50\}$ for Charades-STA and AcitivityNet and $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5,20\}$ for TACoS.
-7.2 MVMR Results
+The existing VMR studies have reported the results of $m\in\{0.3,0.5,0.7\}$ with $n\in\{1,5\}$ (Wang et al., 2022).
+But, since our MVMR setting searches for massive videos, we report the results of $m\in\{0.5,0.7\}$ with $n\in\{1,5,20,50\}$ for Charades-STA and AcitivityNet and $m\in\{0.5,0.7\}$ with $n\in\{1,5,20\}$ for TACoS.
+7.2. MVMR Experimental Results
+Overconfidence for the Misinformation.
+We evaluate various VMR models (i.g., CET, G2L, and MMN) in both VMR and MVMR settings to determine whether their performance significantly degrades when extending the search coverage to a massive video set.
+Figure 6 shows that all the baseline models undergo significant performance degradation in the MVMR task.
+Surprisingly, although CET and G2L show quite strong performance on the VMR task, their performance on the MVMR task notably drops.
+MMN shows a relatively robust performance than CET and G2L, but it still shows significant performance drops.
+They tend to be easily distracted by wrong videos and show overconfidence in the misinformation.
+These outcomes indicate that the current VMR task performance alone is not enough to verify the faithfulness of a VMR model.
+Figure 6.
+VMR vs.
 MVMR Performance.
-We evaluate the performance of CroCs and the baseline models on the introduced MVMR datasets.
-Table 2 shows the experimental results of CroCs and other baselines for our MVMR task.
-Surprisingly, although CET and BET show quite strong performance on the NLVL task, their performance on the MVMR task notably drops.
-For examples, CET and MMN show average 34.5 and 29.7 degradation on R@1(IOU0.5) for all MVMR datasets.
-These outcomes indicate that the current NLVL task performance alone is not enough to verify that a model is well-trained.
-We reveal that CroCs outperforms all other baselines on MVMR task, and this suggests that our model correctly discriminates features of negative and positive moments.
-Experiments show that the models on Charades-STA and TACoS significantly improve performance from the weakly-supervised learning since similar queries exist frequently in these datasets and the possibility of false-negatives is high.
-However, the performance of ActivityNet is attributed to the hard-negative learning since it includes various and complex text queries.
+We report average scores of R1@0.5 and R5@0.5 to reveal the vulnerability of VMR models to misinformation.
+X and Y axes correspond to the models and the average rank score, respectively.
+MVMR Performance.
+Table 2 shows the performance of CroCs and the baseline models on the introduced MVMR datasets.
+We reveal that CroCs outperforms all other baselines on the MVMR task, and this suggests that our model correctly discriminates features of negative and positive moments.
+In the Charades-STA and TACoS datasets, the models show significant improvement in performance from the weakly-supervised potential negative learning.
+These results are attributed to the frequent presence of similar queries in these datasets and the possibility of false negatives.
+On the other hand, the performance of ActivityNet is attributed to the cross-directional hard-negative learning since it includes various and complex text queries.
+Dataset Model R1@0.5 R1@0.7 R5@0.5 R5@0.7 R20@0.5 R20@0.7 R50@0.5 R50@0.7
+$\text{MVMR}_{\text{Charades-STA}}$ CET 2.96 2.02 11.79 6.97 32.21 19.91 50.54 33.88
+BET 2.99 1.94 10.98 7.00 25.11 16.55 37.86 26.08
+G2L 2.78 1.97 10.59 7.36 34.74 21.33 62.72 44.82
+MMN 12.78 8.40 33.29 22.36 57.83 41.71 73.09 56.75
+CroCs 16.31 10.76 39.05 26.99 62.89 46.39 76.08 60.28
+CroCs 17.38 11.44 39.02 27.91 60.98 46.39 76.29 60.76
+$\text{MVMR}_{\text{ActivityNet}}$ CET 0.90 0.56 2.32 1.75 4.73 3.87 8.02 6.56
+BET 1.16 0.76 2.55 2.03 4.64 3.80 7.10 5.88
+G2L 1.32 0.96 6.43 4.55 21.66 14.81 50.07 32.99
+MMN 10.71 6.93 25.04 16.55 44.50 30.62 62.09 45.76
+CroCs 12.68 7.87 28.47 18.16 46.52 31.12 62.12 44.89
+CroCs 15.58 9.51 34.70 22.40 55.58 38.55 70.02 53.20
+$\text{MVMR}_{\text{TACoS}}$ CET 5.79 3.89 18.88 12.17 42.53 23.70 - -
+BET 5.30 2.97 14.84 8.32 39.76 21.51 - -
+G2L 5.82 4.31 21.51 12.29 42.35 22.52 - -
+MMN 8.61 5.60 25.84 15.47 43.89 23.80 - -
+CroCs 10.75 6.76 27.01 15.28 44.38 25.16 - -
+CroCs 10.75 7.30 29.68 16.50 44.77 24.62 - -
+Table 2.
+Experimental Results on MVMR datasets.
+Bolded and under-lined results indicate the $1^{st}$ and $2^{nd}$ best performance, respectively.
+CroCs is the model using only weakly-supervised potential negative learning (§6.1).
+CroCs uses both of two informative sample-weighted learning methods (§6.1 and §6.2).
+We train all the baselines by three trials and report the averaged MVMR scores.
 MVMR Settings with Video Retrieval.
-The practical usage scenario for the MVMR should include video retrieval.
-Therefore, We also conduct MVMR experiments with a video retrieval model, PRMR model Dong et al.
-(2022), shown as Table 3.
-We construct a pipeline to filter top-5 logit videos from our MVMR retrieval pool using the PRMR model and solve MVMR task.
+The practical usage scenario of VMR uses a pipeline integrating video retrieval with video moment retrieval.
+To analyze in depth the excellence of CroCs, we also conduct MVMR experiments with a pipeline including a video retrieval model, PRMR model (Dong et al., 2022).
+The experimental results with video retrieval are shown in Table 3.
+We construct a pipeline to filter top-5 logit videos from our MVMR retrieval pool using the PRMR model and solve the MVMR task.
 We use the publicly deployed PRMR model on Charades-STA and ActivityNet datasets.
 These experiments prove that video retrieval models are helpful to increase MVMR performance, and CroCs still shows superior performance than MMN in this pipeline.
-Model R@1 IOU0.3 R@1 IOU0.5 R@5 IOU0.3 R@5 IOU0.5
-$\text{MMN}_{\text{Charades-STA}}$ 32.15 27.10 65.97 57.66
-$\text{CroCs}_{\text{Charades-STA}}$ 36.64 29.27 68.06 59.78
-$\text{MMN}_{\text{ActivityNet}}$ 32.39 25.04 66.84 55.03
-$\text{CroCs}_{\text{ActivityNet}}$ 39.25 29.59 71.54 60.02
-Table 3: MVMR Experiments with Video Retrieval.
-8 Conclusion
-In this paper, we propose the Massive Videos Moment Retrieval (MVMR) task, which aims to detect a moment for a natural language query from a massive video set.
-To stimulate the research, we propose a simple automatic method to construct reliable MVMR datasets and build three MVMR datasets using existing NLVL datasets.
-We introduce a robust training method called CroCs to distinguish positives from negative distractors effectively in solving MVMR task.
-Limitations
-Our dataset construction framework automatically defines positive and negative sets using semantic similarity evaluation.
-Unlike the existing VCMR framework, which overlooks the risk of labeling errors in its automated dataset construction process, our MVMR framework introduces some filtering methods to mitigate labeling errors.
-However, our method still faces limitations due to its reliance on automatic labeling, which fundamentally carries a risk of labeling inaccuracies.
-To eliminate the risk of labeling errors thoroughly, a constructed dataset should be reviewed manually through complete enumeration.
-Appendix A Datasets Details
-Charades-STA
-is an extended dataset of action recognition and localization dataset Charades Sigurdsson et al.
-(2016) by Gao et al.
-(2017) for video moment retrieval.
-It is comprised of 5,338 videos and 12,408 query-moment pairs in the training set, and 1,334 videos and 3,720 query-moment pairs in the test set.
-We report evaluation results on test set in our experiments.
-ActivityNet-Captions
-is constructed on ActivityNet v1.3 dataset Heilbron et al.
-(2015), where included videos cover various complex human actions.
-It is originally designed for video captioning, and recently used into video moment retrieval.
-It contains 37,417, 17,505, and 17,031 query-moment pairs for training, validation, and testing respectively.
-We report the evaluation result on val_2 set following the setting of the MMN paper Wang et al.
-(2022).
-TACoS
-includes 127 videos selected from the MPII-Cooking dataset.
-It consists of 18,818 query-moment pairs of various cooking activities in the kitchen annotated by Regneri et al.
-(2013).
-It Gao et al.
-(2017) consists of 10,146, 4,589, and 4,083 query-moment pairs for training, validation and testing, respectively.
-We report evaluation results on test set in our experiments.
-Appendix B Analysis on Filter Methods
-We construct reliable MVMR datasets using SimCSE and EMScore.
-The visualizations of the similarity scores derived by SimCSE and EMScore is shown as Figure 7.
-For the SimCSE distribution, we select target queries of "person turn a light on", "The person peels a kiwi", and "He is using a push mower to mow the grass" for the Charades-STA, TACoS, and ActivityNet, respectively.
-Figure 6 shows the distribution of SimCSE and human-annotated STS16Agirre et al.
-(2016) dataset scores.
+Dataset Model R1@0.3 R1@0.5 R5@0.3 R5@0.5
+$\text{MVMR}_{\text{Charades-STA}}$ MMN 32.15 27.10 65.97 57.66
+CroCs 36.64 29.27 68.06 59.78
+$\text{MVMR}_{\text{ActivityNet}}$ MMN 32.39 25.04 66.84 55.03
+CroCs 39.25 29.59 71.54 60.02
+Table 3. MVMR Experiments with Video Retrieval.
+8. Conclusion
+In this paper, we propose the Massive Videos Moment Retrieval for Faithfulness Evaluation (MVMR) task, which aims to detect a moment for a natural language query from a massive video set to evaluate the faithfulness of video moment retrieval.
+To stimulate the research, we propose a simple automatic method to construct trustful MVMR datasets by employing semantic distance verification methods, and finally build three MVMR datasets by extending existing VMR datasets.
+We further introduce a robust informative sample-weighted learning method called CroCs to distinguish positives from negative distractors effectively in solving the MVMR task.
+Acknowledgements.
+This work was supported by Adobe Research and LG AI Research.
+This work was partly supported by Samsung Electronics.
+This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)].
+K.
+Jung is with ASRI, Seoul National University, Korea.
+The Institute of Engineering Research at Seoul National University provided research facilities for this work.
+Appendix A Analysis on Filter Methods
+We construct trustful MVMR datasets using SimCSE and EMScore.
+The visualizations of the results derived by SimCSE and EMScore for the TACoS dataset are shown in Figure 7-(a).
+We select ”The person peels a kiwi” to show the distribution of SimCSE embeddings.
+Figure 7. Qualitative Analysis for Filtering Methods for the TACoS Dataset.
+Figure 8 shows the distribution of SimCSE score and human-annotated STS16(Agirre et al., 2016) dataset score.
 STS16 dataset provides text pairs and annotated similarity scores (five-point Likert scale).
@@
 Consequently, we can reliably discern the similarity between queries using the provided thresholds.
-Figure 6: The distribution of SimCSE and Human-annotated Scores.
+Figure 8.
+The distribution of SimCSE and Human-annotated Scores.
 The X and Y axes mean the human-annotated and the SimCSE scores, respectively.
 The width of each graph corresponds to the number of samples.
-[…]
-[…]
-Figure 7: The Visulization of the SimCSE and EMScore Similarity Distribution.
-Appendix C Implementation Details
+Appendix B Implementation Details
 We use standard off-the-shelf video feature extractors without any fine-tuning.
-We use VGG Simonyan and Zisserman (2015) feature for Charades-STA and C3D Tran et al.
-(2015) feature for ActivityNet and TACoS following the previous work Wang et al.
-(2022).
-We train CroCs and MMN on 1 NVIDIA RTX A6000 GPU and early stop by averaging scores of all evaluation metrics.
-C.1 CroCs
+We use VGG (Simonyan and Zisserman, 2015) feature for Charades-STA and C3D (Tran et al., 2015) feature for ActivityNet and TACoS following the previous work (Wang et al., 2022).
+We train all VMR models on 1 NVIDIA RTX A6000 GPU and early stop by averaging scores of all evaluation metrics.
+B.1. CroCs
 We implement CroCs by following MMN’s model architecture.
@@
 We calculate the representation of each query $f^{q}\in\mathbb{R}^{d}$ using the global average pooling over all tokens.
-We adopt the approach of encoding the input video as a 2D temporal moment feature map, inspired by 2D-TAN Zhang et al.
-(2020b) and MMN.
+We adopt the approach of encoding the input video as a 2D temporal moment feature map, inspired by 2D-TAN (Zhang et al., 2020a) and MMN.
 To achieve this, we segment the input video into video clips denoted as $\{c_{i}\}_{i=1}^{l_{c}/k}$, where each clip $c_{i}$ consists of $k$ frames.
@@
 We set the dimension of the joint feature visual-text space $d=256$, and temperatures $\tau_{m}=\tau_{q}=0.1$.
-We utilize the pre-trained HuggingFace implementation of DistilBERT Wolf et al.
-(2019).
+We utilize the pre-trained HuggingFace implementation of DistilBERT (Wolf et al., 2019).
 We set margin $\psi$ as 0.4, 0.3, and 0.1 for Charades-STA, ActivityNet, and TACoS, respectively.
-We use AdamW Loshchilov and Hutter (2017) optimizer with learning rate of $1\times 10^{-4}$, $8\times 10^{-4}$, $1.5\times 10^{-3}$ for Charades-STA, ActivityNet, and TACoS, respectively.
+We use AdamW (Loshchilov and Hutter, 2017) optimizer with learning rate of $1\times 10^{-4}$, $8\times 10^{-4}$, $1.5\times 10^{-3}$ for Charades-STA, ActivityNet, and TACoS, respectively.
 We set $\lambda$ as 0.05 for Charades-STA and TACoS and 0.1 for ActivityNet.
@@
 The experimental results for various hyper-parameters for the hard-negative sampling are shown in Figure 9.
-Figure 8: Evaluation results for various hyper-parameters.
-We search various hyper-parameters for the hard negative sampling.
-We measure R@1 (IoU>m), where m$\in\{0.3,0.5,0.7\}$ to find the best hyper-parameters.
-In this figure, each $(x,y)$ means the number of negative samples.
+Figure 9.
+Evaluation results for various hyper-parameters.
+We measure Rank 1 (IoU¿m), where m$\in\{0.3,0.5,0.7\}$ to find the best hyper-parameters.
+Each $(x,y)$ means the number of negative samples.
 $x$ and $y$ mean a number of negative samples for $p^{\prime}(v|q)$ (query-to-video) and $p^{\prime}(q|v)$ (query-to-video), respectively.
-Consequently, we select the number of the CM-ANS samples to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
-C.2 Baselines
-MMN
-Wang et al.
-(2022) adopts a dual-modal encoding design to get video clip and text representations.
+We select the number of hard negative samples (§6.2) to calculate $p^{\prime}(q|v)$ as 100, 200, and 100 and $p^{\prime}(v|q)$ as 50, 100, and 5 for Charades-STA, ActivityNet, and TACoS, respectively.
+B.2. Baselines
+MMN (Wang et al., 2022) adopts a dual-modal encoding design to get video clip and text representations.
 We utilized the publicly deployed code [^1] https://github.com/MCG-NJU/MMN to implement our own MMN model for our experiment using exactly the same hyper-parameters setting.
-G2L
-Li et al.
-(2023) uses a semantically aligned and uniform video grounding framework via geodesic and game theory.
-We utilized the publicly deployed code [^3] https://github.com/lihxxxxx/G2L?tab=readme-ov-file to implement our own G2L model for our experiment using exactly the same hyper-parameters setting proposed by the paper.
-Cross-Encoder Transformer (CET)
-is a cross-modal encoding model implemented based on a 6-layered transformer architecture.
+G2L (Li et al., 2023) is a sort of stat-of-the-art model, extending MMN architecture.
+We utilized the publicly deployed code [^2] https://github.com/lihxxxxx/G2L to implement our own G2L model for our experiment using exactly the same hyper-parameters setting.
+Cross-Encoder Transformer (CET) is a cross-modal encoding model implemented based on a 6-layered transformer architecture.
 It is designed to predict the start and end points of a video moment by utilizing a model architecture, which is used in the QA task.
 We concatenate the pre-extracted video clip features and the text features derived from the Distil-BERT to use as an input of the transformer.
-To concatenate two extracted features, we should equalize the dimensions of two features with a linear layer as follows:
-$\bar{H}^{v}=H^{v}W^{v}$ (7)
-$\bar{H}^{q}=H^{q}W^{q}$
-where $H^{v}\in\mathbb{R}^{l_{v},d^{v}}$, $H^{q}\in\mathbb{R}^{l_{q},d^{q}}$, $W^{v}\in\mathbb{R}^{d^{v},d}$, $W^{q}\in\mathbb{R}^{d^{q},d}$.
+To concatenate two extracted features, we should equalize the dimensions of two features with a linear layer as $\bar{H}^{v}=H^{v}W^{v}$ and $\bar{H}^{q}=H^{q}W^{q}$, where $H^{v}\in\mathbb{R}^{l_{v},d^{v}}$, $H^{q}\in\mathbb{R}^{l_{q},d^{q}}$, $W^{v}\in\mathbb{R}^{d^{v},d}$, $W^{q}\in\mathbb{R}^{d^{q},d}$.
 $H^{v}$ is the pre-extracted video clip features and $H^{q}$ is the features extracted using Distil-BERT.
 And then, we derive the representation of each video clip feature from the last layer of the transformer-encoder as follows:
-$H^{v,q}=TransformerEnc(\bar{H}^{v}\oplus\bar{H}^{q})$ (8)
+(7) $H^{v,q}=TransformerEnc(\bar{H}^{v}\oplus\bar{H}^{q})$
 where $H^{v,q}\in\mathbb{R}^{l_{vq},d}$, $l_{vq}=l_{v}+l_{q}$.
 We utilize only the representation part of video features, and the 3-layered MLPs and Sigmoid follow them to predict the start and end positions of the moment, respectively, as follows:
-$p^{s}=Sigmoid(MLP_{s}(H^{v,q}_{1:l_{v}}))$ (9)
-$p^{e}=Sigmoid(MLP_{e}(H^{v,q}_{1:l_{v}}))$
+(8) $p^{s}=Sigmoid(MLP_{s}(H^{v,q}_{1:l_{v}})),\>\>\>\ p^{e}=Sigmoid(MLP_{e}(H^{v,q}_{1:l_{v}}))$
 where $p^{s}$ and $p^{e}\in\mathbb{R}^{l_{v}}$. We also introduce the interpolation labels $y_{i}^{s}$, $y_{i}^{e}$ to utilize abundant information of moment labels as follows:
-$y_{i}^{s}=\left\{\begin{array}[]{ c l }({\frac{x_{e}-i}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>\>$ (10)
-$y_{i}^{e}=\left\{\begin{array}[]{ c l }({\frac{i-x_{s}}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.\>\>\>\>\>$
-We finally use the binary cross-entropy to calculate a loss for the pair of $(p_{i}^{s},y_{i}^{s})$ and $(p_{i}^{e},y_{i}^{e})$ as follows:
-$\mathcal{L}^{s}=BCE(p^{s},y^{s})$ (11)
-$\mathcal{L}^{e}=BCE(p^{e},y^{e})$
-where BCE means the binary cross-entropy. The final loss to train the CET model is defined as follows:
-$\mathcal{L}=\frac{\mathcal{L}^{s}+\mathcal{L}^{e}}{2}$ (12)
+(9) $y_{i}^{s}=\left\{\begin{array}[]{ c l }({\frac{x_{e}-i}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.$
+$y_{i}^{e}=\left\{\begin{array}[]{ c l }({\frac{i-x_{s}}{x_{e}^{*}-x_{s}^{*}+1}})^{k}&\quad\textrm{if }x_{s}^{*}\leq i\leq x_{e}^{*}\\ 0&\quad\textrm{otherwise}\end{array}\right.$
+We finally use the binary cross-entropy to calculate losses for the pair of $(p_{i}^{s},y_{i}^{s})$ and $(p_{i}^{e},y_{i}^{e})$ as $\mathcal{L}^{s}=BCE(p^{s},y^{s})$ and $\mathcal{L}^{e}=BCE(p^{e},y^{e})$, respectively.
+BCE means the binary cross-entropy.
+The final loss to train the CET model is defined as follows:
+(10) $\mathcal{L}=\frac{\mathcal{L}^{s}+\mathcal{L}^{e}}{2}$
 We use $k=10$ and $d=512$ as a hyperparameter and AdamW optimizer with learning rate 1e-3.
-Bi-Encoder Transformer (BET)
-is a dual-modal encoding model that encodes text and video clip features independently.
+Bi-Encoder Transformer (BET) is a dual-modal encoding model that encodes text and video clip features independently.
 The pre-extracted video clip features are encoded using a 6-layered transformer, and text features are encoded using the Distil-BERT, followed by the average-pooling function and 3-layered MLPs.
-$\tilde{H}^{v}=TransformerEnc(\bar{H}^{v})$ (13)
-$\tilde{H}^{s}=MLP_{s}(Pool(\bar{H}^{q}))$ (14)
-$\tilde{H}^{e}=MLP_{e}(Pool(\bar{H}^{q}))$
+(11) $\tilde{H}^{v}=TransformerEnc(\bar{H}^{v})$
+(12) $\tilde{H}^{s}=MLP_{s}(Pool(\bar{H}^{q})),\>\>\>\ \tilde{H}^{e}=MLP_{e}(Pool(\bar{H}^{q}))$
 where $\tilde{H}^{v}\in\mathbb{R}^{d}$, $\tilde{H}^{s}\in\mathbb{R}^{d}$ and $\tilde{H}^{e}\in\mathbb{R}^{d}$.
@@
 We calculate the cosine similarity between the encoded video clip representations and the encoded start and end representations to predict a moment.
-$p^{s}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{s}))$ (15)
-$p^{e}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{e}))$
+(13) $p^{s}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{s})),\>\>\>\ p^{e}=Sigmoid(r*cosine(\tilde{H}^{v},\tilde{H}^{e}))$
 We use the interpolation labels and the binary cross-entropy loss for the BET, similar to CET. We use $r=10$ and $d=512$ as a hyperparameter and AdamW optimizer with learning rate 1e-3.
-Appendix D Examples of False Negative Videos
-Figure 9: Examples of constructed MVMR datasets.
-We visualize positive and negative video sets for the query “person closes the door" of three constructed datasets.
-A green solid box means a golden positive moment, and blue solid boxes show moments assigned to videos classified as positive.
-The underlined queries mean the most similar query described in Section 4.1 (max aggregation).
-Illustrative instances of constructed MVMR datasets can be found in Figure 9.
-This figure shows video instances that have been identified as similar to a specific target query.
-The target queries correspond to queries of Charades-STA, ActivityNet, and TACoS datasets, sequentially.
-Appendix E Human Evaluation for the constructed MVMR Datasets
-Figure 10: Human Evaluation Sheet Example.
-We conduct a human evaluation on our constructed MVMR datasets to reveal that our method construct MVMR datasets effectively.
-We recruit crowd workers fluent in English through the university’s online community.
-The recruited annotators were provided with a detailed description of task definitions, instructions.
-An example of a sheet we use to conduct a human evaluation is shown in Figure 10.